### v2

# A Unified Framework for N-Dimensional Visualization and Simulation: Implementation and Evaluation including 4D Boolean Operations

Abstract
This study proposes a unified framework for simulation and visualization of intuitive exploration of phenomena in $N$-dimensional space. While specialized libraries offer powerful geometric algorithms, they typically lack integrated environments for interactive trial and error, creating a barrier for researchers. The contribution of this research is the integration of Quickhull-based mesh generation, visualization via hyperplane slicing, and computationally expensive Boolean operations into a single, extensible platform, while maintaining interactivity. To validate its effectiveness, this paper presents a 4-dimensional implementation and introduces a new interaction design, termed ‘High-Dimensional FPS,’ to enable intuitive high-dimensional exploration. Furthermore, as a case study to demonstrate the framework’s high extensibility, I also integrated a non-rigid body physics simulation based on Extended Position Based Dynamics (XPBD). Experimental results confirmed the effectiveness of the proposed method, achieving real-time rendering (80 fps) of complex 4D objects and completing Boolean operations within seconds in a standard PC environment. By providing an accessible and interactive platform, this work lowers the entry barrier for high-dimensional simulation research and enhances its potential for applications in education and entertainment.
1 Introduction
While the simulation and visualization of high-dimensional spaces are theoretically established to some extent, practical integrated environments and concrete methods are limited. This study proposes a unified framework for $N$-dimensional simulation and visualization that can be easily adopted and applied by researchers and developers. This framework unifies previously disparate techniques—such as Quickhull-based high-dimensional convex hull mesh generation, visualization based on hyperplane slicing, high-dimensional exploration through view transformations, and mesh processing via Boolean operations—enabling their integrated use and providing an environment for the intuitive understanding of high dimensions.
Previously, functionalities such as high-dimensional mesh generation and Boolean operations were offered within specialized computational geometry libraries like the Computational Geometry Algorithms Library (CGAL). However, an environment that combined these with visualization and physics simulation to enable interactive trial and error has been lacking.
The framework proposed in this paper is positioned as one solution to this challenge. Its architecture, founded on the design principle of separating topology and geometry, achieves high extensibility. I demonstrate its effectiveness through a case study integrating a non-rigid body physics simulation based on Extended Position Based Dynamics (XPBD), showing that external components can be easily integrated. This approach is expected to lower the barrier to entry for high-dimensional simulation research, enhance its potential for applications in education and entertainment, and contribute to the future growth of the research community and academic advancement.
2 Related Work
Prior research in high-dimensional simulation and visualization has primarily focused on reporting individual techniques. In mesh generation, the construction of convex hull meshes using the Quickhull algorithm is widely known [quickhull], but it has a high computational cost to guarantee mathematical rigor. In physics simulation, while simulations of non-rigid bodies based on XPBD have been reported for 3D space [XPBD], and research on the simulation of rigid bodies in 4D and higher dimensions exists [nrigid], there are exceedingly few reports of integrated environments that allow for the interactive manipulation of non-rigid bodies and the visualization of objects in 4D or higher dimensions. Turning to mesh editing, Boolean operations are a standard technique in 3D. In higher-dimensional spaces, specialized computational geometry libraries like CGAL also provide this functionality as an offline computation. However, these processes require significant computational resources to guarantee mathematical precision and are not intended for interactive use that involves real-time result verification. Regarding visualization, while a GPU-based 4D visualization architecture has been proposed [GL4D], its primary focus is on fast rendering, and it offers a limited environment for the kind of exploratory experience that this paper aims to achieve.
3 Proposed Method
3.1 Design Philosophy and Architectural Overview
This section describes the technical details of the proposed unified framework. This framework integrates the functionalities of mesh generation, Boolean operations, visualization, and physics simulation into a single, extensibility-focused pipeline. I will first provide an overview of the entire architecture, and then detail the core algorithms that constitute each function and their implementation.
In the design and implementation of this framework, three principles were emphasized: algorithmic transparency, minimal reliance on external libraries, and future extensibility. This is a design decision that prioritizes the framework’s utility as a research platform—allowing researchers to easily understand, verify, and incorporate new functionalities into the high-dimensional geometry algorithms—over extreme performance optimization. Based on this philosophy, all computational geometry logic is unified in a standard C# implementation on the CPU. While GPU-based parallelization can deliver high performance for specific tasks, it introduces issues of code complexity and hardware dependency. The CPU-centric approach of this paper aims to enhance platform independence and algorithmic transparency. This implementation strategy is expected to lower the technical barriers for other researchers to verify this study’s algorithms or apply them to their own research. Furthermore, within the scope of this paper, numerical robustness related to floating-point arithmetic is excluded from consideration and is left for future work. The software for this research was implemented in C# on the Unity engine and is available as an open-source project [Arai_2025_UnifiedFramework].
3.2 Overview of the Overall Architecture
Figure 1: An overview of the proposed framework. The left side shows the static shape definition workflow, and the right side shows the dynamic visualization pipeline.
As illustrated in Figure 1, the overall architecture and data flow of this framework consist of multiple loosely coupled components: convex hull mesh generation, mesh editing, and visualization. The mesh generation component uses a Quickhull-based algorithm to construct an index list of facets from the input point cloud data, which defines the topology of the $N$-dimensional object. The mesh editing component dynamically processes and modifies the mesh’s topology and geometry based on user operations, such as Boolean operations. The visualization component generates and renders a 3D cross-section each frame by performing hyperplane slicing on the coordinate-transformed mesh, based on the current vertex coordinates and the user’s viewpoint.
Each component interacts only through a well-defined data structure for the object’s topology and geometry, thereby ensuring high modularity. Furthermore, I have adopted Geometric Algebra as a unified mathematical foundation that enables the consistent handling of geometric entities such as $N$-dimensional vectors, planes, and rotations. For Geometric Algebra itself, which I use as my foundation, [vince2008] is cited as an excellent introductory text for researchers in computer graphics.
3.3 N-Dimensional Convex Hull Construction
Figure 2: Stages of vertex management and convex hull construction in Direct Quickhull. Orange vertices are unprocessed, gray vertices are processed, gray facets are interior, and black facets form the outer hull. In the iterative step, the convex hull is expanded using a selected facet from the hull (red) and its farthest point (red).
3.3.1 Comparison with Existing Methods
This component is centered around a self-contained C# implementation of the classic $N$-dimensional Quickhull algorithm [quickhull]. A standard implementation would perform an optimization called candidate point pruning to maximize computational efficiency. This strategy, in which each facet individually manages a list of candidate points and excludes points absorbed into the convex hull’s interior from computation, is extremely effective for volumetric point clouds containing numerous interior points. However, the primary applications within my framework—such as the basic convex hull mesh construction for initial shapes and dynamic simplicial tessellation (detailed in Section 3.3.3)—involve frequently handling small-scale, surface-like point clouds, typically with fewer than several dozen vertices and almost no interior points. In such contexts, the performance gain offered by a pruning mechanism is limited, while the added implementation complexity reduces code maintainability and introduces risks to robustness in a floating-point arithmetic environment.
Therefore, my implementation considers this trade-off and intentionally omits the pruning mechanism. As shown in Figure 2, my algorithm performs a linear search over a single, global set of unprocessed points to find the farthest point, and removes only those points that are used as vertices of the convex hull from the set. This simplified state-management approach, in exchange for theoretical computational efficiency on large volumetric point clouds, provides the practically significant advantages of implementation simplicity and independence from external libraries for my primary use cases. Hereafter in this paper, I will refer to this method as ”Direct Quickhull.”
3.3.2 Calculating N-Dimensional Normals and Visible Facets
In Direct Quickhull, efficiently computing facet normal vectors plays a crucial role in determining overall performance. The conventional approach based on null-space computation is unsuitable for real-time applications, as it typically requires methods such as Singular Value Decomposition (SVD). To solve this problem, my implementation introduces the concepts of the wedge product and the dual from Geometric Algebra. First, in an $N$-dimensional space $\mathbb{R}^{N}$, I define a facet from its NNvertices $\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{N-1}\}$and construct the (N1)(N-1) linearly independent vectors that span the facet.
$\mathbf{v}_{i}=\mathbf{p}_{i}-\mathbf{p}_{0}\quad(\text{for }i=1,\dots,N-1)$ (1)
Next, I calculate the wedge product of these vectors.
$\mathbf{B}=\mathbf{v}_{1}\wedge\mathbf{v}_{2}\wedge\dots\wedge\mathbf{v}_{N-1}$ (2)
[…]
$\mathbf{n}=\mathbf{B}^{*}$ (3)
This series of operations can ultimately be implemented efficiently as the calculation of the determinant of an $n\times n$ matrix, as follows:
$\mathbf{n}:=\det\begin{vmatrix}\mathbf{e}_{1}&\mathbf{e}_{2}&\dots&\mathbf{e}_{n}\\ v_{1,1}&v_{1,2}&\dots&v_{1,n}\\ v_{2,1}&v_{2,2}&\dots&v_{2,n}\\ \vdots&\vdots&\ddots&\vdots\\ v_{n-1,1}&v_{n-1,2}&\dots&v_{n-1,n}\end{vmatrix}$ (4)
Next, it is necessary to correctly determine the orientation of the facet’s normal vector. In conventional 3D, the correct orientation of a normal vector for a facet on the surface of a convex hull was obtained by computing the vertex winding order, which relies on the right-handed rule of the cross product. However, the cross product is unique to 3D space. In higher dimensions, there are no geometric concepts analogous to the winding order that intuitively define a face’s orientation.
To solve this problem, my implementation uses the geometric property that the normal vectors of a convex hull’s facets must always point outward. This orientation is determined by calculating the dot product of the facet’s normal vector and a vector pointing from a point on the facet to the centroid of the initial simplex constructed at the beginning of the algorithm. If the sign of the dot product is negative, the normal vector is determined to be pointing outward. If the sign is positive, its sign is inverted, thereby ensuring that I always obtain a normal vector with the correct outward-facing orientation.
Figure 3: The logic for determining visible facets in Quickhull. This illustrates the method for identifying facets that are visible from the farthest point $\mathbf{P}_{\text{new}}$. The centroid of the facet being tested is denoted as $\mathbf{P}_{\text{c}}$, and its normal vector is $\mathbf{n}$. The vector pointing from $\mathbf{P}_{\text{new}}$ to $\mathbf{P}_{\text{c}}$ is defined as $\mathbf{v}$.
Next, I use the facet normal vectors to determine the visible facets, which is a necessary step in the process of expanding the convex hull.
As shown in Figure 3, the facets visible from $\mathbf{P_{\text{new}}}$ can be found by examining the sign of the dot product of $\mathbf{v}$ and $\mathbf{n}$. The convex hull can then be expanded by connecting the new vertex $\mathbf{P_{\text{new}}}$ to the boundary of the identified set of visible facets (i.e., the set of ridges that are not shared among the visible facets).
3.3.3 Method for Simplicial Tessellation during Convex Hull Construction
my implementation not only computes the facets that constitute the surface of the convex hull but is also equipped with the functionality to simultaneously generate its interior volume as a set of $N$-dimensional simplices. This is accomplished by applying the incremental process of the Quickhull algorithm. In Quickhull’s incremental process, the newly added volume is represented as a set of $N$-dimensional simplices formed from the new vertex $\mathbf{p}_{\text{new}}$ and the visible facets. my implementation directly uses this principle, appending the corresponding simplices to a list. By repeating this process, I simultaneously obtain both the surface facets of the convex hull and a list of simplices that perfectly tessellates its interior volume. This simplicial decomposition data serves two important roles within this framework. Its primary application is to serve as the base data for rapidly performing a simplicial decomposition of the interior of facets clipped during Boolean operations. Furthermore, it is intended for future application in efficiently determining the cross-sectional structure of facets from hyperplane slicing in spaces of five or more dimensions, a topic detailed in Section 3.5.3.
3.4 N-Dimensional Boolean Operations
3.4.1 Algorithm Overview and Design Philosophy
While Boolean operations are an extremely powerful method for interactively modifying the shape of an $N$-dimensional mesh, their interactive implementation in high-dimensional space presents a significant challenge from a computational cost perspective. To address this challenge, my framework implements a practical four-stage algorithm. First, in a broad phase, I use spatial partitioning to narrow down potential intersection candidates. Next, in a narrow phase, I determine the actual geometric intersections. The intersected facets are then decomposed into smaller facets in a tessellation stage. Finally, the results are classified in an inside-outside test stage to construct the final operation result.
In designing this algorithm, I prioritized a balance of three factors over extreme performance optimization or a full guarantee of mathematical rigor: implementation simplicity, extensibility to $N$-dimensions, and a practical level of performance that enables interactive trial and error. The following sections will detail the specific techniques used in each stage and the design decisions behind them.
3.4.2 Broad Phase: Candidate Pruning via Spatial Partitioning
In $N$-dimensional Boolean operations, a brute-force intersection test between all facets of two meshes, with $N_{\text{A}}$ and $N_{\text{B}}$ facets respectively, has a computational complexity of $O(N_{\text{A}}N_{\text{B}})$, making it a major bottleneck in interactive implementations. To avoid this problem, my implementation introduces a spatial partitioning grid as the broad phase to efficiently prune pairs of facets that could potentially intersect. I compute the axis-aligned bounding box (AABB) for each facet and register the facet with the grid cells that its AABB occupies. As a result, only facet pairs that share a common cell become candidates for the precise intersection tests in the subsequent narrow phase. Compared to hierarchical spatial partitioning methods such as BVH or Octree, the uniform grid approach adopted in my implementation offers the advantage of being simple to implement while still achieving practical pruning performance for my target use cases.
Figure 4: The sequential tessellation process of a facet in a Boolean operation (analogized in 3D). After the input facet (yellow) is clipped and tessellated by a facet from the opposing object (blue, left), each resulting facet is further tessellated by a subsequent blue facet, as shown in the center. By repeating this process for all intersecting facets, a mesh (Output) that is accurately partitioned at the boundary with the opposing object is ultimately generated.
3.4.3 Narrow Phase: Facet-Facet Intersection Test
In the narrow phase, I perform precise geometric intersection tests on the candidate pairs pruned during the broad phase. The principle of this algorithm is a generalization to $N$-dimensions of the method proposed by Tomas Möller et al. for fast triangle-triangle intersection tests in 3D [Moller1997]. Its essence lies in decomposing the intersection test between two facets, $\mathcal{F}_{\text{A}}$ and $\mathcal{F}_{\text{B}}$, into two symmetric sub-problems. First, I compute the geometric intersection region, $\mathcal{I}_{\text{B}}$, between the (N1)(N-1)-dimensional hyperplane spanned by $\mathcal{F}_{\text{A}}$ and the facet $\mathcal{F}_{\text{B}}$. Second, I perform the same operation with the roles swapped, computing the intersection region $\mathcal{I}_{\text{A}}$ between the hyperplane spanned by $\mathcal{I}_{\text{B}}$ and the facet $\mathcal{F}_{\text{A}}$. Both of these intersection regions are (N2)(N-2)-dimensional convex polytopes. Finally, I determine if an intersection between the facets exists by testing whether $\mathcal{I}_{\text{A}}$ and $\mathcal{I}_{\text{B}}$ overlap within their common (N2)(N-2)-dimensional space.
If an intersection is found to exist, this information is utilized in the subsequent tessellation stage. Specifically, I register this intersection pair in a dictionary where one facet is the key and the other is the value. By performing this process for all intersection pairs, I can systematically construct information about the clipping relationships, enabling us to handle complex situations where one facet intersects with multiple other facets.
In my 4D implementation, the intersection region is a 2D convex polygon. Since the number of vertices is small, I adopted the Separating Axis Theorem (SAT), which is relatively straightforward to implement. However, in five or more dimensions, the maximum number of vertices of the (N2)(N-2)-dimensional polytope intersection region increases, making the implementation of SAT more complex. Therefore, for extension to higher dimensions, an intersection test approach using the GJK algorithm is considered a more practical option.
3.4.4 Dynamic Tessellation via Direct Quickhull
A facet that is found to intersect in the narrow phase is clipped by the (N1)(N-1)-dimensional hyperplane spanned by the opposing facet, and the clipped facet is then tessellated into multiple smaller facets. This entire process is illustrated in Figure 4 using an analogy in 3D. As the figure shows, clipping can decompose the original facet, an (N1)(N-1)-dimensional simplex, into an (N1)(N-1)-dimensional convex polytope with more than NNvertices. Consequently, it is necessary to efficiently re-tessellate the resulting polytope back into multiple smaller facets.
The core of this tessellation algorithm is to classify the vertices of the facet into the positive side (P-side) and the negative side (N-side) based on their position relative to the clipping hyperplane. The tessellation process is executed only when the facet has vertices in both half-spaces; otherwise, no tessellation is performed.
However, in an implementation that assumes floating-point arithmetic, this theoretical classification alone is insufficient. Problems can arise from numerical errors, such as an intersection that should theoretically be a single point being detected as multiple points, or the classification becoming unstable when a vertex is extremely close to the hyperplane. To address this challenge, my implementation introduces a two-stage vertex merging process. First, it unifies points that are in close proximity to each other within the computed set of intersection vertices. Second, it evaluates the distance between each vertex of the merged intersection region and the vertices of the original facet; if they are close, the intersection vertex is completely snapped to the coordinates of the facet vertex. While this operation may introduce minor deformations to the mesh geometry, it plays a crucial role in maintaining the topological consistency of the original facet within the tessellated mesh.
After the merging process, the final tessellation is performed. First, I apply the simplicial tessellation feature of Direct Quickhull to the combined surface-like point cloud of the P-side vertices and the intersection region vertices to tessellate the P-side region into multiple smaller facets. Next, a similar operation is performed for the N-side vertices and the intersection region vertices to tessellate the N-side region. Through these two processes, the original facet is tessellated such that its mesh structure is consistent with the clipping plane. This approach demonstrates that Direct Quickhull can effectively use its capabilities even in dynamic mesh processing, thereby guaranteeing the $N$-dimensional extensibility of the entire algorithm.
3.4.5 Inside-Outside Testing and Result Construction
Figure 5: Classification of general and degenerate intersection cases. This figure categorizes the general case, where a ray crosses a single facet, and representative degenerate cases that require precise intersection tests, where the ray passes exactly through a facet, edge, or vertex.
To construct the final result of a Boolean operation—namely, the union, intersection, or difference from the set of tessellated facets, it is necessary to classify whether each facet is inside or outside the opposing object. For this inside-outside test, my implementation adopts the ray casting method. The fundamental principle of this method is based on intersection parity. A ray is cast from the centroid of the facet under test toward the opposing surface, and the number of intersections is counted. The state is then determined according to whether this number is even or odd, and the same logic applies in $N$-dimensional space. In my implementation, the ray is parallel to the x-axis, and its endpoint is set to a point slightly beyond the opposing mesh’s AABB. This ensures both computational efficiency and stability. To make this method operate robustly in a floating-point environment, it is necessary to handle degenerate cases where the ray passes through a facet’s vertices or edges, as shown in Figure 6. In an implementation that introduces a tolerance, an intersection with a vertex or edge, which should theoretically be a single point of intersection, can be detected as multiple, slightly different intersection points with the various facets that share that feature. This disrupts the parity of the intersection count and causes the inside-outside test to fail. To prevent this misclassification, my implementation employs a two-stage solution. First, the coordinates of the detected intersection points are discretized based on a tolerance, and multiple proximate intersection points are grouped into a single ’intersection event’ so they can be handled uniformly in subsequent processing. Next, the nature of this grouped intersection event is determined. Specifically, it is necessary to distinguish whether the ray is piercing the surface or merely touching it.
Figure 6: An example of a degenerate case in intersection testing. In the figure, $\mathbf{P_{1}}$, $\mathbf{P_{2}}$, and $\mathbf{P_{3}}$ are the vertices of a 2D polygon; $\mathbf{n_{1}}$ and $\mathbf{n_{2}}$ are the outward-facing normal vectors of the edges connected to vertex $\mathbf{P_{2}}$; and $\mathbf{R_{t}}$ and $\mathbf{R_{a}}$ are two different rays passing through $\mathbf{P_{2}}$. The left diagram shows a piercing case, and the right diagram shows a tangent case.
To do this, I evaluate the dot product of the ray’s direction vector and the normal vector of every facet included in the event. If the signs of the dot products are biased to one side (either all positive or all negative), the event is considered to be piercing the surface, and the intersection count is incremented by one. If, on the other hand, the signs are distributed between both positive and negative, the ray is determined to be just tangent to the surface, and it is not included in the intersection count. Based on this inside-outside test executed through this series of processes, the tessellated facets are classified according to the type of Boolean operation, and the final mesh is constructed.
However, in cases where the ray is coplanar with a facet, such as the ”Co-planar Facet” case shown in Figure 5, the dot product of the ray and the normal vector is theoretically zero, causing the sign-based test to fail robustly. Therefore, there is no guarantee that intersections can always be correctly determined under these conditions. A practical approach would be to avoid such degenerate geometric configurations by using existing techniques, such as ray perturbation.
3.5 Visualization of N-Dimensional Objects
3.5.1 Cross-Sectioning as the Basic Strategy
This component is responsible for generating and rendering 3D cross-sections based on $N$-dimensional mesh data. The basic strategy to achieve this transformation is based on the concept of sequentially reducing the dimension. This is a process of recursively clipping an $N$-dimensional object with an (N1)(N-1)-dimensional hyperplane, and then clipping the resulting cross-section with an (N2)(N-2)-dimensional hyperplane. In this paper, I refer to this method of sequentially reducing dimensions as Hierarchical Cross-Sectioning. With this approach, the complex problem of computing cross-sections in arbitrary dimensions can always be reduced to the simpler problem of the intersection between an $N$-dimensional facet and an (N1)(N-1)-dimensional hyperplane. In the following sections, as a concrete application of this Hierarchical Cross-Sectioning for $N=4$, I will detail the specific algorithm for generating 3D cross-sections from 4D objects and discuss its implementation challenges.
3.5.2 Facet Clipping Algorithm in the 4D Implementation
To generate a 3D cross-section of an object in 4D space, my implementation computes the geometric intersection between each facet of the 4D object—a tetrahedron—and an axis-parallel hyperplane $w=w_{\text{slice}}$ (where $w=w_{\text{slice}}$ is an arbitrary constant). This procedure is executed independently for every facet of the input mesh according to the following steps.
First, I determine on which side of the clipping hyperplane each of the facet’s four vertices is located. If all vertices are on the same side, the facet does not intersect the hyperplane and skip intersection computation. If the vertices straddle the hyperplane, the intersection point of each constituent edge of the facet with the hyperplane is calculated via linear interpolation. Let the two vertices defining an edge be
$\mathbf{p_{0}}=({x_{0}},{y_{0}},{z_{0}},{w_{0}})\quad\mathbf{p_{1}}=({x_{1}},{y_{1}},{z_{1}},{w_{1}})$ (5)
The interpolation parameter ttfor determining the position of the intersection point is then calculated from the $w$-coordinates of both vertices and the $w$-coordinate of the clipping plane, $w_{\text{slice}}$, as shown in Equation (6).
$t=\frac{w_{\text{slice}}-w_{0}}{w_{1}-w_{0}}$ (6)
Next, using this parameter $t$, the coordinates of the intersection point $\mathbf{p}_{\text{isect}}$ in 4D space are calculated according to Equation (7).
$\mathbf{p}_{\text{isect}}=\mathbf{p}_{1}+t(\mathbf{p}_{1}-\mathbf{p}_{0})$ (7)
The 3D coordinates, obtained by extracting the $x,y,z$ components of the resulting 4D coordinate vector $\mathbf{p}_{\text{isect}}$, directly become the coordinates of a vertex that constitutes the cross-section.
By performing this operation for all edges that intersect the hyperplane, I obtain the set of vertices of the cross-section, which are then connected to form a polygon. The shape of the cross-section generated by the intersection of a single 4D facet and a 3D hyperplane is, within 3D space, either a triangle with three vertices or a convex polygon with four vertices. A 4-vertex convex polygon can be divided into two triangles to obtain a set of triangular polygons for rendering.
3.5.3 Simplicial Decomposition of the Cross-Section
Figure 7: Dimensional dependence of the number of cross-section vertices generated during hyperplane clipping. When clipping a facet in 3D, the number of vertices is always two. However, when clipping a 3D facet in 4D, the number varies between three and four depending on the relative position of the hyperplane, and is thus not uniquely determined. Regarding the color coding in this figure, the blue plane represents the clipping hyperplane, and the red area indicates the portion of it that exists on the dimension being clipped. The darker shades are highlights for visual clarity.
The cross-section generated by the above clipping algorithm is generally an (N2)(N-2)-dimensional convex polytope. For subsequent processing, this must be decomposed into a set of (N2)(N-2)-dimensional simplices. In the 4D implementation, the cross-section is a 2D convex polygon (either a triangle or a quadrilateral) with a maximum of four vertices, as shown in Figure 7. Although the order of the vertex list is undefined, the polygon can be safely tessellate into two triangles by sorting the vertices by angle around the polygon’s centroid and then systematically partitioning it. my implementation efficiently performs this sort using only vector operations, avoiding computationally expensive trigonometric functions.
However, this angle-sorting method cannot be generalized to five or more dimensions $(N\geq 5)$. The maximum number of vertices of an (N2)(N-2)-dimensional cross-section, obtained by clipping an $N$-dimensional facet (which has NNvertices) with a hyperplane, is given by the well-known result in combinatorics shown in Equation (8).
$\left\lfloor\frac{N}{2}\right\rfloor\cdot\left\lceil\frac{N}{2}\right\rceil$ (8)
For example, in 5D, the cross-section can be a 3D convex polytope with a maximum of $\lfloor 5/2\rfloor\cdot\lceil 5/2\rceil=6$ vertices. The simplicial tessellation of such a complex polytope is a non-trivial computational problem. As a dimension-independent and general solution to this high-dimensional extensibility challenge, I can use the simplicial tessellation feature of Direct Quickhull, which was detailed in Section 3.3.3. By providing the set of vertices that constitute the cross-section as input to Direct Quickhull, it is possible to efficiently tessellate the interior of the cross-sectional polytope itself into a set of (N1)(N-1)-dimensional facets. As mentioned in Section 3.3.1, my implementation, which omits the interior point search, is well-suited for processing this type of surface-like point cloud. This ensures that the visualization pipeline of my framework is, in principle, extensible to higher dimensions.
3.5.4 Robust Method for Determining Cross-Section Normals
When generating a 3D cross-section for rendering from a 4D object, it is necessary to determine the winding order of the vertices that form the 3D facet. The clipping algorithm from the previous section produces a list of vertices in an undefined order. If this unordered vertex list is used as is, a serious problem arises: many rendering pipelines determine the front and back of a face based on the input order, which would result in inconsistent face orientations and prevent the correct rendering of the shape.
To solve this problem, my method utilizes the normal vector of the original 4D facet as a reference for the correct orientation. Specifically, I first project the original 4D normal vector into 3D space by extracting its $x,y,z$ components, and define this as the true normal vector that the 3D facets should follow. Next, I compute a temporary normal vector from the generated vertex set using the cross product and evaluate its dot product with the true normal vector. If the sign of this dot product is negative, the winding order of the vertices is reversed. This ensures that all triangles constituting the cross-sectional mesh have a consistent vertex winding order that can be correctly interpreted by a general-purpose renderer.
3.6 Pose Representation and Control Scheme for Interactive Exploration
3.6.1 Pose Representation and Management with Geometric Algebra Rotors
In representing the pose of an $N$-dimensional object, conventional 3D methods like Euler angles or quaternions cannot be directly applied due to issues such as gimbal lock and difficulties in extending them to higher dimensions. To solve this challenge, my framework adheres to the mathematical foundation of Geometric Algebra, as mentioned in Section 3.2, and represents the object’s pose uniformly using a Rotor.
A rotor is a rotation operator defined by a specific plane of rotation and a rotation angle, which allows the composition of rotations and their application to vectors to be handled as an algebraic product, regardless of the dimension. A bivector $\mathbf{B}$ representing the plane of rotation is parallel to the plane spanned by two vectors, $\mathbf{a}$ and $\mathbf{b}$, and can be defined by their wedge product:
$\mathbf{B}=\mathbf{a}\wedge\mathbf{b}$ (9)
A rotor is an operator that represents a rotation of an arbitrary angle with respect to this plane. The rotor $R$ is defined using the normalized bivector $\mathbf{\hat{B}}$ as follows:
$R=\cos\frac{\theta}{2}+\sin\frac{\theta}{2}\mathbf{\hat{B}}$ (10)
The rotation of a vector is then expressed using the rotor as shown in the following equation:
$\mathbf{v}^{\prime}=R\mathbf{v}\tilde{R}$ (11)
Here, $\tilde{R}$ is the reverse of the rotor $R$. If $R$ is normalized, $\tilde{R}$ is equivalent to $R^{-1}$.
In my implementation, instead of having the user manipulate the rotor directly, I allow them to specify the amount of rotation for each of the six fundamental rotation planes in 4D space $(xy,xz,xw,yz,yw,zw)$ as individual parameters, similar to Euler angles. These parameter values are then composed in a predefined order to internally construct the final rotation operator, the rotor. This design, which separates the intuitive parameters provided to the user from the mathematically robust operator used internally, completely avoids gimbal lock and is easily extensible to higher dimensions. Furthermore, my implementation fixes the clipping hyperplane to be axis-parallel $(w=\text{const})$ and transforms the object’s pose instead. From the user’s perspective, this approach is visually equivalent to clipping a stationary object with an arbitrarily oriented hyperplane. However, it replaces the computationally expensive problem of arbitrary hyperplane clipping with a combination of an efficient pose transformation and an axis-parallel clip, enabling interactive observation of the cross-sections.
3.6.2 Camera-Based Coordinate Transformation
This section defines the view transformation in $N$-dimensions, which transforms vertices from the world coordinate system to the camera coordinate system based on the camera’s state (position and pose).
Figure 8: A conceptual diagram of the coordinate transformation from the world coordinate system to the camera coordinate system. Let the camera’s position be $\mathbf{C}$ and the target point be $\mathbf{P}$. The relative position vector of the target point, $\mathbf{v}$, is defined as the difference between $\mathbf{P}$ and $\mathbf{C}$. The components obtained by projecting $\mathbf{v}$ onto the basis vectors that represent the camera’s orientation, $\mathbf{e}_{\text{right}}$ and $\mathbf{e}_{\text{fwd}}$, directly become the coordinate values of $\mathbf{P}$ in the camera coordinate system. In this setup, the camera’s position $\mathbf{C}$ corresponds to the origin of the camera coordinate system.
To intuitively understand the geometric essence of this transformation, I consider the 2D space shown in Figure 8 as an example. I denote the position of a target vertex $\mathbf{P}$ in the world coordinate system as $\mathbf{P}^{(\mathrm{W})}$ and the position of the camera $\mathbf{C}$ as $\mathbf{C}^{(\mathrm{W})}$. The coordinate transformation takes $\mathbf{P}^{(\mathrm{W})}$, $\mathbf{C}^{(\mathrm{W})}$, and the rotor $R$ representing the pose as input. The transformation process first derives the basis vectors that span the camera coordinate system, $\mathbf{e}_{\text{right}}^{(\mathrm{W})}$ and $\mathbf{e}_{\text{fwd}}^{(\mathrm{W})}$, from the rotor $R$ representing the camera’s pose, as shown in Equation (11). Next, the relative position vector $\mathbf{v}^{(\mathrm{W})}$ from the camera’s position to the object’s vertex is calculated as shown in Equation (12).
$\mathbf{v}^{(\mathrm{W})}=\mathbf{P}^{(\mathrm{W})}-\mathbf{C}^{(\mathrm{W})}$ (12)
Finally, this relative vector $\mathbf{v}^{(\mathrm{W})}$ is projected onto each basis vector via the dot product, and the resulting projection lengths determine the respective components of $\mathbf{P}^{(\mathrm{C})}$.
$\mathbf{P}^{(\mathrm{C})}_{x}$ $=\mathbf{v}^{(\mathrm{W})}\cdot\mathbf{e}_{\text{right}}^{(\mathrm{W})}$ (13)
$\mathbf{P}^{(\mathrm{C})}_{y}$ $=\mathbf{v}^{(\mathrm{W})}\cdot\mathbf{e}_{\text{fwd}}^{(\mathrm{W})}$
This entire sequence of processes can be interpreted as two separate geometric operations. Equation (12) corresponds to a translation that aligns the origin of the world coordinate system with the camera’s position. Equation (13) corresponds to a rotation that aligns the coordinate system with the camera’s orientation. By using homogeneous coordinates, these two operations can be combined into a single $(N+1)\times(N+1)$ matrix, analogous to the view matrix commonly used in computer graphics. The description using dot products in this paper is specifically intended to directly illustrate the geometric meaning of the rotation operation.
3.6.3 Design of the Control Scheme for Interactive Exploration
To address the challenge of mapping the six independent rotational degrees of freedom required to control the camera’s orientation in 4D space onto a standard mouse and keyboard, I adopted two design principles: orthogonality of controls and leveraging existing control conventions (specifically, those of First-Person Shooters, or FPS). The 3D translations (WASD keys) and pitch rotation (mouse vertical movement) follow standard conventions. The primary challenge, which is the assignment of a rotation to the horizontal mouse movement ($x$-axis), is resolved by a modal input switch using a modifier key (Left Control). In its normal mode, horizontal mouse movement controls the primary yaw rotation ($xz$-plane) for 3D exploration. While the modifier key is held down, this input is switched to control the rotation in the $xw$-plane, which serves a similar role in 4D exploration. This design ensures that a single input always corresponds orthogonally to a single operation, resulting in a highly predictable control scheme that allows the user to operate while being clearly conscious of switching between 3D and 4D exploration modes.
3.6.4 Extension of the Manipulation System to N Dimensions
This design principle can be generalized to $N$-dimensions. I apply a systematic approach to address the rotational degrees of freedom in an $N$-dimensional space, which increase quadratically as N(N1)/2N(N-1)/2. First, translations and rotations within the 3D cross-section that the user directly observes are kept on the standard control scheme (WASD/mouse), regardless of the total number of dimensions. This provides the user with a consistently stable operational base. Next, for the remaining (N(N1)/2)3(N(N-1)/2)-3 rotations outside the observed space, I consider an approach that generalizes the design of my 4D implementation. Specifically, horizontal mouse movement is consistently assigned to rotations that occur outside the observed space; the user selects the target high-dimensional axis $(w,u,\ldots)$ via a combination of modifier keys, which in turn manipulates the corresponding rotation plane $(xw,xu,\ldots)$. Because this approach dedicates the primary rotational input to high-dimensional exploration while allowing the control scheme to be systematically extended, it is expected to provide a scalable foundation for exploring $N$-dimensional space while preserving control orthogonality.
3.7 Demonstration of Extensibility
The core of this framework’s extensibility lies in the design principle of separating topology and geometry. This section demonstrates the effectiveness of this design through a case study in which a 4D non-rigid body physics simulation is integrated.
A practical challenge in this integration is the mismatch of data formats. For computational efficiency, the physics simulation requires unique, non-duplicated vertex data. The rendering pipeline, on the other hand, requires duplicated vertex data to accommodate correct normals and UVs. The separated architecture of my framework resolves this discrepancy through index mapping. During initialization, the physics engine extracts the set of unique vertices from the rendering mesh’s geometry and builds a map, just once, that maintains the correspondence between the indices of the unique vertices and the indices of the original, duplicated vertices. During the simulation, computations are performed only on the set of unique vertices. Then, in each frame, this map is used to efficiently update the entire vertex array for rendering.
This pattern—maintaining an internally optimized data format and converting it to the required format only during communication is a general solution for integrating any external component that requires a different data format, and is not limited to this case study. This demonstrates that the design principle of separating topology and geometry is what enables the high modularity and extensibility of my framework.
4 Plex: An N-Dimensional Mesh Data Exchange Ecosystem
Table 1: Overview of standard chunks in the .plex format
Chunk Type Requirement Description
’META’ Required Stores metadata for the entire mesh, such as the number of dimensions and vertices.
’VERT’/’VERD’ Required Stores the geometry data of vertex coordinates in single or double precision.
’FACE’ Required Stores the vertex indices (topology) that constitute the facets.
’NORM’/’NORMD’ Required Stores the normal vectors that define the orientation of the facets in single or double precision.
’CENT’/’CENTD’ Optional Stores the cached centroid coordinates of facets to reduce computational load.
Table 2: Internal data layout of the ’META’ chunk
Offset Size Type Description
0x00 4 bytes int32 Dimension: N
0x04 4 bytes uint32 Reserved: Padded with 0.
0x08 8 bytes uint64 Vertex Count: V
0x10 8 bytes uint64 Creation Time: Unix Time Stamp
0x18 8 bytes uint64 Modified Time: Unix Time Stamp
0x20 1 byte uint8 Precision Flag: 0=single precision, 1=double precision
0x21 3 bytes byte[3] Reserved: Padded with 0.
0x24 4 bytes uint32 LSoftware: Byte length of the following software name.
0x28 Variable char[] Software Name: UTF-8 encoded string.
The $N$-dimensional simulation environments targeted by this research require, as their foundation, an extensible data exchange ecosystem for uniformly handling high-dimensional mesh data. However, to date, no widely accepted standard format for exchanging $N$-dimensional mesh data has existed, leading to a fragmented situation where formats vary among researchers and projects. This state of affairs has been a barrier to the smooth sharing of data between different tools and research groups. To address this challenge, this study defines a systematic data exchange ecosystem for $N$-dimensional meshes, named ”Plex” (with the .plex extension), which is a file format featuring a chunk-based architecture.
4.1 Dual-Format Strategy
Research workflows that handle $N$-dimensional data present two distinct requirements: the need to inspect and verify data content in the early stages, and the need for high processing performance in large-scale computations. To meet these requirements, Plex adopts a dual-format strategy, following the design philosophy proven by the glTF format in the 3D domain.
The first format is a schema based on the standard JSON technology, which prioritizes data readability. This allows for the easy inspection and debugging of data content using a rich ecosystem of existing tools. The second is a chunk-based binary format, .plex, which prioritizes execution performance. This format is intended for fast loading by applications and efficient data storage.
The reference implementation provided in this study includes a one-way converter from the .plex format to the JSON format. This enables researchers to typically work with the high-performance binary format while exporting to a human-readable format as needed, thereby achieving a balance between execution efficiency and data transparency.
4.2 .plex Binary Format Specification
The .plex format is designed with a chunk-based architecture, following the proven extensibility model of standard formats such as PNG. A file consists of a global header followed by a sequence of at least four self-descriptive chunks as shown in Table 1; the internal structure of the META chunk is detailed in Table 2. Each chunk contains a 4-character chunk type, the data length, the data payload, and a CRC32 checksum to verify data integrity. The data payload structure for the primary chunks, excluding the META chunk, is defined as follows. The payload of the FACE chunk, which defines the topology, first specifies the total number of facets as a 64-bit unsigned integer, followed by a list of vertex indices (an array of 32-bit unsigned integers) that constitute each facet. In contrast, the payload for the VERT or VERD chunks, which store vertex coordinates, contains no metadata such as counts; for a mesh with $V$ vertices in $N$ dimensions, it consists solely of a contiguous array of $V\times N$ numerical values. The respective data types are 32-bit single-precision and 64-bit double-precision floating-point numbers. The payload layout for the NORM/NRMD chunks (storing facet normals) and the optional CENT/CNTD chunks (storing centroids) also conforms to this VERT/VERD chunk format.
A parser compliant with the specification is guaranteed to safely skip any unknown chunk types it cannot interpret. This ensures that backward compatibility with existing software is maintained even if the community adds new data in the future, such as textures or animations, as custom chunks.
Furthermore, to encourage community-driven extensions, a naming convention is established that reserves identifiers containing one or more lowercase letters for use as custom chunks. This standardization will facilitate data sharing among researchers and serve as a technical foundation for improving the reproducibility of published results. Moreover, the forward compatibility guaranteed by this specification allows for future extensions, providing a sustainable infrastructure that transforms $N$-dimensional mesh data from something previously confined to individual environments into a resource that can be shared and reused by the entire community.
Figure 9: The set of base objects used in the experiments. The top row shows the 3D objects, and the bottom row shows the 4D objects. The number to the right of each object’s name is the number of facets that constitute its mesh. The 4D objects are visualized using the hyperplane slicing feature of my framework, with their 3D cross-sections at three representative $w$-coordinates (w=0.8,0,0.8)(w=-0.8,0,0.8) displayed side-by-side.
5 Experiments and Evaluation
5.1 Experimental Setup
The performance evaluation in this section is not intended to be a speed comparison with existing offline methods. Instead, its purpose is to evaluate the performance scalability of the proposed CPU-based framework with respect to the scale of the processed mesh (e.g., the number of facets). All implementation and performance evaluations in this study were conducted in the following environment.
  • Hardware Environment:
    • CPU: Intel Core i7-12700H (2.30 GHz, 14Cores)
    • Memory: 16 $\mathrm{GB}$ RAM
    • GPU: Intel Iris Xe Graphics (Integrated, Shared Memory)
  • Software Environment:
    • OS: Windows 11 Home (64-bit)
    • Development Environment: Unity 2022.3 LTS (2022.3.62f2)
    • Programming Language: C#
5.2 Algorithmic Performance Evaluation
Table 3: Performance evaluation of the Direct Quickhull implementation. This table shows the average processing time for convex hull construction with randomly placed vertices (vertex count: 10 to 1000) in 3D and 4D spaces. The processing time is the average of 100 trials.
Dimension Vertices Time ($\mathrm{ms}$) Avg. Facets
3D 10 $\centering 0.05\@add@centering$ $\centering 13.52\@add@centering$
3D 20 $\centering 0.12\@add@centering$ $\centering 25.32\@add@centering$
3D 50 $\centering 0.35\@add@centering$ $\centering 48.20\@add@centering$
3D 100 $\centering 0.88\@add@centering$ $\centering 73.34\@add@centering$
3D 200 $\centering 2.50\@add@centering$ $\centering 110.66\@add@centering$
3D 500 $\centering 9.43\@add@centering$ $\centering 185.24\@add@centering$
3D 1000 $\centering 26.95\@add@centering$ $\centering 268.32\@add@centering$
4D 10 $\centering 0.21\@add@centering$ $\centering 20.38\@add@centering$
4D 20 $\centering 0.56\@add@centering$ $\centering 43.11\@add@centering$
4D 50 $\centering 1.64\@add@centering$ $\centering 100.02\@add@centering$
4D 100 $\centering 3.84\@add@centering$ $\centering 174.32\@add@centering$
4D 200 $\centering 7.97\@add@centering$ $\centering 275.43\@add@centering$
4D 500 $\centering 27.43\@add@centering$ $\centering 496.94\@add@centering$
4D 1000 $\centering 75.99\@add@centering$ $\centering 779.91\@add@centering$
Table 4: Performance evaluation results for Boolean operations. I performed union, intersection, and difference operations in 3D and 4D spaces using combinations of the base objects shown in Figure 9. For each combination, the table shows the processing time and the size (number of facets) of the resulting mesh. The processing time is the average of three trials.
Input Union Intersection Difference
Dimension ObjA ObjB Time ($\mathrm{ms}$) Size Time ($\mathrm{ms}$) Size Time ($\mathrm{ms}$) Size
3D Cube Cube $\centering 15.42\@add@centering$ $\centering 50\@add@centering$ $\centering 13.47\@add@centering$ $\centering 26\@add@centering$ $\centering 14.47\@add@centering$ $\centering 38\@add@centering$
3D Cube Pole $\centering 11.55\@add@centering$ $\centering 56\@add@centering$ $\centering 10.62\@add@centering$ $\centering 24\@add@centering$ $\centering 12.43\@add@centering$ $\centering 36\@add@centering$
3D Cube Random $\centering 20.87\@add@centering$ $\centering 87\@add@centering$ $\centering 18.60\@add@centering$ $\centering 59\@add@centering$ $\centering 20.46\@add@centering$ $\centering 79\@add@centering$
3D Cube Hollow Cube $\centering 14.49\@add@centering$ $\centering 104\@add@centering$ $\centering 11.82\@add@centering$ $\centering 28\@add@centering$ $\centering 14.65\@add@centering$ $\centering 88\@add@centering$
3D Pole Pole $\centering 13.75\@add@centering$ $\centering 64\@add@centering$ $\centering 11.46\@add@centering$ $\centering 32\@add@centering$ $\centering 10.95\@add@centering$ $\centering 48\@add@centering$
3D Pole Random $\centering 25.39\@add@centering$ $\centering 86\@add@centering$ $\centering 20.78\@add@centering$ $\centering 38\@add@centering$ $\centering 19.72\@add@centering$ $\centering 74\@add@centering$
3D Pole Hollow Cube $\centering 16.90\@add@centering$ $\centering 156\@add@centering$ $\centering 12.86\@add@centering$ $\centering 84\@add@centering$ $\centering 18.29\@add@centering$ $\centering 144\@add@centering$
3D Random Random $\centering 33.65\@add@centering$ $\centering 98\@add@centering$ $\centering 28.24\@add@centering$ $\centering 54\@add@centering$ $\centering 24.50\@add@centering$ $\centering 78\@add@centering$
3D Random Hollow Cube $\centering 21.56\@add@centering$ $\centering 126\@add@centering$ $\centering 20.94\@add@centering$ $\centering 56\@add@centering$ $\centering 17.82\@add@centering$ $\centering 94\@add@centering$
3D Hollow Cube Hollow Cube $\centering 18.93\@add@centering$ $\centering 170\@add@centering$ $\centering 14.92\@add@centering$ $\centering 66\@add@centering$ $\centering 14.97\@add@centering$ $\centering 118\@add@centering$
4D Cube Cube $\centering 323.44\@add@centering$ $\centering 1023\@add@centering$ $\centering 291.94\@add@centering$ $\centering 301\@add@centering$ $\centering 333.86\@add@centering$ $\centering 694\@add@centering$
4D Cube Pole $\centering 205.06\@add@centering$ $\centering 768\@add@centering$ $\centering 186.23\@add@centering$ $\centering 299\@add@centering$ $\centering 176.74\@add@centering$ $\centering 324\@add@centering$
4D Cube Random $\centering 2091.09\@add@centering$ $\centering 6142\@add@centering$ $\centering 2081.24\@add@centering$ $\centering 3777\@add@centering$ $\centering 2079.20\@add@centering$ $\centering 4143\@add@centering$
4D Cube Hollow Cube $\centering 682.97\@add@centering$ $\centering 3017\@add@centering$ $\centering 656.85\@add@centering$ $\centering 1471\@add@centering$ $\centering 661.75\@add@centering$ $\centering 2693\@add@centering$
4D Pole Pole $\centering 139.65\@add@centering$ $\centering 857\@add@centering$ $\centering 128.59\@add@centering$ $\centering 480\@add@centering$ $\centering 129.56\@add@centering$ $\centering 631\@add@centering$
4D Pole Random $\centering 1225.35\@add@centering$ $\centering 2095\@add@centering$ $\centering 1213.80\@add@centering$ $\centering 1123\@add@centering$ $\centering 1237.35\@add@centering$ $\centering 1655\@add@centering$
4D Pole Hollow Cube $\centering 638.94\@add@centering$ $\centering 3357\@add@centering$ $\centering 626.98\@add@centering$ $\centering 1353\@add@centering$ $\centering 639.27\@add@centering$ $\centering 3402\@add@centering$
4D Random Random $\centering 2175.25\@add@centering$ $\centering 4449\@add@centering$ $\centering 2126.16\@add@centering$ $\centering 2326\@add@centering$ $\centering 2157.43\@add@centering$ $\centering 3114\@add@centering$
4D Random Hollow Cube $\centering 2406.69\@add@centering$ $\centering 9633\@add@centering$ $\centering 2332.84\@add@centering$ $\centering 5292\@add@centering$ $\centering 2418.27\@add@centering$ $\centering 9181\@add@centering$
4D Hollow Cube Hollow Cube $\centering 1598.95\@add@centering$ $\centering 7973\@add@centering$ $\centering 1559.87\@add@centering$ $\centering 4763\@add@centering$ $\centering 1576.44\@add@centering$ $\centering 6801\@add@centering$
Table 5: Interactive performance evaluation. This table shows the rendering performance in scenes containing 1 to 20 4D objects (hypercubes with 47 facets each), while the user continuously performed movement and view rotation operations in the xzxz and xwxw planes for 30 seconds. Each value is the average of three 30 second measurement trials.
Objects (Total Facets) Average FPS Minimum FPS 99th Percentile Frametime ($\mathrm{ms}$)
1(47) 77.9 29.8 19.9
5(235) 58.3 27.9 23.4
10(470) 38.6 16.8 32.5
15(705) 27.0 14.4 46.9
20(940) 21.2 13.1 58.1
5.2.1 Performance of the Direct Quickhull Implementation
As shown in Table 3, for small-scale point clouds with 50 or fewer vertices which are the intended primary use case in my framework the processing time is under 1.7 $\mathrm{ms}$ in either dimension, confirming that its performance is sufficient for this purpose.
5.2.2 Performance of N-Dimensional Boolean Operations
Next, I evaluated the performance of the $N$-dimensional Boolean operations, which are central to interactive mesh editing. I executed union, intersection, and difference operations between various objects in 3D and 4D spaces and measured their computation times. The results are presented in Table 4. In 3D space, the operations were generally completed in several tens of milliseconds, depending on the combination of objects and the type of operation. In 4D space, however, the computational load increased significantly. Even a simple combination of two Cubes required approximately 300 $\mathrm{ms}$, and in some cases, operations between two complex Random objects took over two seconds to complete. These results indicate that high-dimensional Boolean operations can be executed on a time scale of a few seconds. This processing time is considered to be within an acceptable range for an interactive workflow, where the user waits for the result of an operation before considering the next action. This demonstrates the feasibility of incorporating these operations, which were previously dominated by offline batch processing, into interactive applications.
5.3 Interactive Performance Evaluation
This section quantitatively evaluates the performance characteristics of the proposed framework. I placed multiple simple 4D objects, each with 47 facets, and measured various performance metrics as the total number of facets in the scene was increased. The experimental results are shown in Table 5. These results indicate that the performance of my framework scales with the total number of facets in the scene. Furthermore, a specific performance limit was identified: the ability to maintain 30 fps, a common benchmark for interactive operation, is limited to a total facet count of approximately 500 (equivalent to about 10 objects). This measurement provides a performance baseline for evaluating the effects of future improvements, such as a GPU implementation.
6 Conclusion and Future Work
In this paper, I proposed an interactive framework for the unified simulation and visualization of $N$-dimensional space. The core of this framework lies in the design principle of separating topology and geometry. This allowed for the integration of functionalities such as Quickhull-based mesh generation, hyperplane slicing, and $N$-dimensional Boolean operations onto a single platform. The effectiveness of this design was demonstrated through the integration of an XPBD-based physics simulation. I also proposed a control scheme for high-dimensional exploration and the Plex data exchange format.
The proposed algorithm can be extended to arbitrary dimensions; however, in this study, the evaluation was performed on the 4-dimensional implementation. Extension to five or more dimensions may introduce new performance bottlenecks due to the increased combinatorial complexity in intersection testing and tessellation. Furthermore, major future challenges include the development of a more robust $N$-dimensional non-rigid body model, application to new phenomena such as fluid simulations, and a GPU implementation of the parallelizable components currently handled by the CPU.
Acknowledgments
I would like to express my sincere gratitude to Mr. Hiroaki Naito of Waseda Osaka High School for his constant and enthusiastic guidance throughout this research. I also wish to express my deepest appreciation to Professor Yasushi Homma of Waseda University for his valuable professional advice on the mathematical expressions related to Geometric Algebra used in this paper. Finally, I would like to thank the anonymous reviewers for their constructive comments, which helped to improve this manuscript.

### v3

@@
-A Unified Framework for N-Dimensional Visualization and Simulation: Implementation and Evaluation including 4D Boolean Operations
+A Unified Architecture for N-Dimensional Visualization and Simulation: 4D Implementation and Evaluation including Boolean Operations
 Abstract
-This study proposes a unified framework for simulation and visualization of intuitive exploration of phenomena in $N$-dimensional space.
-While specialized libraries offer powerful geometric algorithms, they typically lack integrated environments for interactive trial and error, creating a barrier for researchers.
-The contribution of this research is the integration of Quickhull-based mesh generation, visualization via hyperplane slicing, and computationally expensive Boolean operations into a single, extensible platform, while maintaining interactivity.
-To validate its effectiveness, this paper presents a 4-dimensional implementation and introduces a new interaction design, termed ‘High-Dimensional FPS,’ to enable intuitive high-dimensional exploration.
-Furthermore, as a case study to demonstrate the framework’s high extensibility, I also integrated a non-rigid body physics simulation based on Extended Position Based Dynamics (XPBD).
-Experimental results confirmed the effectiveness of the proposed method, achieving real-time rendering (80 fps) of complex 4D objects and completing Boolean operations within seconds in a standard PC environment.
-By providing an accessible and interactive platform, this work lowers the entry barrier for high-dimensional simulation research and enhances its potential for applications in education and entertainment.
+This paper proposes a unified software architecture for visualization and simulation based on a design targeting an $N$-dimensional space.
+The contribution of this work lies in presenting an architectural configuration that integrates multiple processes into a single software architecture: Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (including orientation and view transformations), and hyperplane slicing for visualization.
+The proposed approach adopts an approximate implementation that tolerates numerical errors and prioritizes implementation transparency over guarantees of numerical rigor.
+The experimental results and evaluations presented in this paper are limited to a 4D implementation; no evaluation is conducted for $N>4$, and the discussion is restricted to stating that the architecture itself has a dimension-independent structure.
+This paper also proposes an interaction design for high-dimensional exploration based on FPS navigation.
+As an input example involving shape changes over time, a non-rigid body simulation based on XPBD (Extended Position Based Dynamics) is integrated into the 4D implementation.
+Experimental results confirm that the 4D implementation runs on a single PC.
 1 Introduction
-While the simulation and visualization of high-dimensional spaces are theoretically established to some extent, practical integrated environments and concrete methods are limited.
-This study proposes a unified framework for $N$-dimensional simulation and visualization that can be easily adopted and applied by researchers and developers.
-This framework unifies previously disparate techniques—such as Quickhull-based high-dimensional convex hull mesh generation, visualization based on hyperplane slicing, high-dimensional exploration through view transformations, and mesh processing via Boolean operations—enabling their integrated use and providing an environment for the intuitive understanding of high dimensions.
-Previously, functionalities such as high-dimensional mesh generation and Boolean operations were offered within specialized computational geometry libraries like the Computational Geometry Algorithms Library (CGAL).
-However, an environment that combined these with visualization and physics simulation to enable interactive trial and error has been lacking.
-The framework proposed in this paper is positioned as one solution to this challenge.
-Its architecture, founded on the design principle of separating topology and geometry, achieves high extensibility.
-I demonstrate its effectiveness through a case study integrating a non-rigid body physics simulation based on Extended Position Based Dynamics (XPBD), showing that external components can be easily integrated.
-This approach is expected to lower the barrier to entry for high-dimensional simulation research, enhance its potential for applications in education and entertainment, and contribute to the future growth of the research community and academic advancement.
+This paper builds on prior work on visualization and simulation in high-dimensional spaces.
+However, examples remain limited that integrate Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (orientation and view transformations), and hyperplane slicing for visualization into a single software architecture, under the assumption of an iterative workflow of user actions and result inspection.
+In this work, we propose a unified architecture intended for use by researchers and developers.
+While this architecture is designed for an $N$-dimensional space, the present study implements and evaluates it in 4D.
+The contribution of this work is to present an architectural configuration that integrates the above processing components as a single software architecture and to evaluate the 4D implementation experimentally.
+In this paper, interactivity is defined as a workflow that presupposes repeated cycles of user actions and result inspection, in contrast to offline batch processing in which input conditions are fixed in advance and executed in a single run.
+Hereafter, the term “interactive,” including its adjectival and adverbial usages, refers to usage scenarios that satisfy or require this notion of interactivity.
 2 Related Work
-Prior research in high-dimensional simulation and visualization has primarily focused on reporting individual techniques.
-In mesh generation, the construction of convex hull meshes using the Quickhull algorithm is widely known [quickhull], but it has a high computational cost to guarantee mathematical rigor.
-In physics simulation, while simulations of non-rigid bodies based on XPBD have been reported for 3D space [XPBD], and research on the simulation of rigid bodies in 4D and higher dimensions exists [nrigid], there are exceedingly few reports of integrated environments that allow for the interactive manipulation of non-rigid bodies and the visualization of objects in 4D or higher dimensions.
-Turning to mesh editing, Boolean operations are a standard technique in 3D.
-In higher-dimensional spaces, specialized computational geometry libraries like CGAL also provide this functionality as an offline computation.
-However, these processes require significant computational resources to guarantee mathematical precision and are not intended for interactive use that involves real-time result verification.
-Regarding visualization, while a GPU-based 4D visualization architecture has been proposed [GL4D], its primary focus is on fast rendering, and it offers a limited environment for the kind of exploratory experience that this paper aims to achieve.
+Prior work on visualization and simulation in high-dimensional spaces is often reported in terms of individual elemental techniques.
+For convex hull computation applicable to high dimensions, the pivoting method [2] and Quickhull [3] are known as representative approaches.
+In implementations that ensure robustness against inputs with numerical errors or degenerate configurations, these algorithms can incur increased computational cost due to exact predicates and additional handling.
+Convex hull computation is part of the fundamental problem set in high-dimensional computational geometry, and textbooks systematically organize problems such as convex hulls, halfspace intersection, and arrangements [18].
+As existing methods for high-dimensional visualization, there are approaches based on projection to lower-dimensional spaces.
+Cook et al.
+provide an overview of methods and design issues for exploring and presenting high-dimensional information via 2D projections [6].
+Dimensional stacking has also been surveyed as a class of visualization techniques that assigns each dimension to a nested 2D layout [12].
+These works mainly survey visualization techniques.
+In contrast, on the implementation side, architectures and frameworks for high-dimensional visualization have also been proposed.
+For mesh-based approaches, a GPU-based 4D visualization architecture has been proposed [5] and discussed from the viewpoint of rendering and visualization of 4D shapes.
+As a software framework for computations on convex polyhedra, “polymake” has been reported [10].
+It provides visualization, exemplified by Schlegel diagrams, together with the underlying algorithmic processing.
+However, these prior examples differ in scope from the unified environment targeted in this paper, which aims to integrate not only visualization but also other processing in an interactive manner.
+For Boolean operations on 3D meshes, methods based on intersection resolution and spatial partitioning have been reported; Zhou et al.
+present a unified framework that handles union, intersection, and difference based on cell decomposition and extraction using winding number vectors [26].
+In the 3D setting, implementation examples of interactive Boolean operations have also been reported [4].
+CGAL adopts a design policy that emphasizes geometric robustness [8].
+Under this policy, CGAL and related libraries provide set operations in 2D and 3D, exemplified by Nef polyhedra, whereas implementation reports of Boolean operations on explicit boundary meshes in 4D and higher dimensions are limited.
+For set operations in 4D and higher dimensions, libraries for high-dimensional computation based on polyhedral representations have been proposed [24].
+However, these assume representations based on halfspace constraints or cell decompositions, rather than explicit boundary meshes.
+They also typically assume offline computation with exact arithmetic, and implementation reports aimed at interactive operating environments remain limited.
+In physical simulation, non-rigid body simulation based on XPBD has been reported in 3D space [15], and there are also studies that handle rigid bodies in 4D and higher dimensions [22].
+However, reports whose primary goal is to visualize non-rigid objects in 4D and higher dimensions and handle them interactively remain limited.
 3 Proposed Method
+3.1 Overview of the Overall Architecture
+Figure 1: An overview of the proposed architecture. The left side shows the static shape definition workflow, and the right side shows the dynamic visualization pipeline.
+This section describes the technical details of the proposed architecture.
+We first provide an overview of the overall architecture, and then describe in detail the algorithms that constitute each function and their implementations.
+As shown in Fig.
+1, the overall structure and data flow of the architecture consist of multiple components: convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing.
+The convex hull mesh generation component employs a Quickhull-based algorithm to construct a list of facet indices that define the topology of an $N$-dimensional object from the input point set data.
+The Boolean operation component dynamically modifies the topology and geometry of the mesh based on user interactions.
+The coordinate transformation component converts vertex coordinates into a camera-referenced coordinate system by applying a view transformation based on the viewpoint, after incorporating orientation transformations updated through user operations.
+The hyperplane slicing component applies hyperplane slicing to the output mesh from the coordinate transformation component and generates and renders 3D cross-sections within the update loop.
+In this study, geometric algebra is adopted as the mathematical foundation for uniformly describing geometric entities such as $N$-dimensional vectors, planes, and rotations.
+Introductory explanations of geometric algebra and its use in computer science and computer graphics can be found in [23], systematic treatments in [7], and survey-style overviews in [14].
 3.2 Design Philosophy and Architectural Overview
-This section describes the technical details of the proposed unified framework.
-This framework integrates the functionalities of mesh generation, Boolean operations, visualization, and physics simulation into a single, extensibility-focused pipeline.
-I will first provide an overview of the entire architecture, and then detail the core algorithms that constitute each function and their implementation.
-In the design and implementation of this framework, three principles were emphasized: algorithmic transparency, minimal reliance on external libraries, and future extensibility.
-This is a design decision that prioritizes the framework’s utility as a research platform—allowing researchers to easily understand, verify, and incorporate new functionalities into the high-dimensional geometry algorithms—over extreme performance optimization.
-Based on this philosophy, all computational geometry logic is unified in a standard C# implementation on the CPU.
+The design policy of this architecture emphasizes three principles: algorithmic transparency, minimization of dependence on external libraries, and separation of topology and geometry.
+This policy is based on a design decision that prioritizes simplification of description and implementation over performance optimization, so that researchers and developers can trace the structure of the algorithms.
+Based on this design policy, all computational geometry logic is unified under a CPU-based implementation written in standard C#.
 While GPU-based parallelization can deliver high performance for specific tasks, it introduces issues of code complexity and hardware dependency.
-The CPU-centric approach of this paper aims to enhance platform independence and algorithmic transparency.
-This implementation strategy is expected to lower the technical barriers for other researchers to verify this study’s algorithms or apply them to their own research.
-Furthermore, within the scope of this paper, numerical robustness related to floating-point arithmetic is excluded from consideration and is left for future work.
-The software for this research was implemented in C# on the Unity engine and is available as an open-source project [Arai_2025_UnifiedFramework].
-3.2 Overview of the Overall Architecture
-Figure 1: An overview of the proposed framework. The left side shows the static shape definition workflow, and the right side shows the dynamic visualization pipeline.
-As illustrated in Figure 1, the overall architecture and data flow of this framework consist of multiple loosely coupled components: convex hull mesh generation, mesh editing, and visualization.
-The mesh generation component uses a Quickhull-based algorithm to construct an index list of facets from the input point cloud data, which defines the topology of the $N$-dimensional object.
-The mesh editing component dynamically processes and modifies the mesh’s topology and geometry based on user operations, such as Boolean operations.
-The visualization component generates and renders a 3D cross-section each frame by performing hyperplane slicing on the coordinate-transformed mesh, based on the current vertex coordinates and the user’s viewpoint.
-Each component interacts only through a well-defined data structure for the object’s topology and geometry, thereby ensuring high modularity.
-Furthermore, I have adopted Geometric Algebra as a unified mathematical foundation that enables the consistent handling of geometric entities such as $N$-dimensional vectors, planes, and rotations.
-For Geometric Algebra itself, which I use as my foundation, [vince2008] is cited as an excellent introductory text for researchers in computer graphics.
-3.3 N-Dimensional Convex Hull Construction
+The CPU-based C# implementation is intended to enhance platform independence and algorithmic transparency.
+In addition, this paper adopts an approximate implementation that tolerates numerical errors and does not guarantee numerical robustness arising from floating-point arithmetic, which is positioned as a subject for future work.
+The 4D implementation in this study was developed in C# on the Unity engine and is released as open-source software [1].
+In this architecture, a design is adopted in which the vertex index sequences that define facets (topology) and the vertex coordinate arrays (geometry) are managed independently.
+This design assumes scenarios in which time-varying geometry is provided as input for a fixed topology.
+Details are described in Section 3.8.
+3.3 N-Dimensional Convex Hull Mesh Generation
 Figure 2: Stages of vertex management and convex hull construction in Direct Quickhull.
@@
 3.3.1 Comparison with Existing Methods
-This component is centered around a self-contained C# implementation of the classic $N$-dimensional Quickhull algorithm [quickhull].
+This component is centered on a self-contained C# implementation of the classical $N$-dimensional Quickhull algorithm [3].
 A standard implementation would perform an optimization called candidate point pruning to maximize computational efficiency.
-This strategy, in which each facet individually manages a list of candidate points and excludes points absorbed into the convex hull’s interior from computation, is extremely effective for volumetric point clouds containing numerous interior points.
-However, the primary applications within my framework—such as the basic convex hull mesh construction for initial shapes and dynamic simplicial tessellation (detailed in Section 3.3.3)—involve frequently handling small-scale, surface-like point clouds, typically with fewer than several dozen vertices and almost no interior points.
-In such contexts, the performance gain offered by a pruning mechanism is limited, while the added implementation complexity reduces code maintainability and introduces risks to robustness in a floating-point arithmetic environment.
-Therefore, my implementation considers this trade-off and intentionally omits the pruning mechanism.
-As shown in Figure 2, my algorithm performs a linear search over a single, global set of unprocessed points to find the farthest point, and removes only those points that are used as vertices of the convex hull from the set.
-This simplified state-management approach, in exchange for theoretical computational efficiency on large volumetric point clouds, provides the practically significant advantages of implementation simplicity and independence from external libraries for my primary use cases.
-Hereafter in this paper, I will refer to this method as ”Direct Quickhull.”
+In this strategy, each facet maintains its own list of candidate points, and points that have been absorbed into the interior of the convex hull are excluded from further computation; this is effective for volume-like point sets containing many interior points.
+However, in this architecture, the primary use cases are basic convex hull mesh construction for initial shapes and dynamic simplex tessellation (described in Section 3.3.3), and it often handles small surface-like point sets with only a few dozen vertices and few interior points.
+In this context, the performance gains from pruning are limited, whereas increased implementation complexity entails a risk of reduced code maintainability.
+Therefore, this architecture deliberately removes the pruning mechanism in light of this trade-off.
+As shown in Fig.
+2, it linearly searches a single global set of candidate points to find the farthest point, and removes from the candidate set only those points that are used as convex hull vertices.
+In this paper, we refer to this approach as “Direct Quickhull.”
 3.3.2 Calculating N-Dimensional Normals and Visible Facets
-In Direct Quickhull, efficiently computing facet normal vectors plays a crucial role in determining overall performance.
-The conventional approach based on null-space computation is unsuitable for real-time applications, as it typically requires methods such as Singular Value Decomposition (SVD).
-To solve this problem, my implementation introduces the concepts of the wedge product and the dual from Geometric Algebra.
-First, in an $N$-dimensional space $\mathbb{R}^{N}$, I define a facet from its NNvertices $\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{N-1}\}$and construct the (N1)(N-1) linearly independent vectors that span the facet.
+In Direct Quickhull, facet normal vectors are computed iteratively, and this computation can become a dominant factor in the overall processing time.
+One possible approach for computing normals is to use null-space computation; however, this typically requires singular value decomposition (SVD) or similar methods, which can impose a substantial computational burden in the iterative calculations assumed for interactive use.
+To address this issue, this architecture introduces the concepts of the wedge product and the dual from geometric Algebra.
+First, in an $N$-dimensional space $\mathbb{R}^{N}$, we define a facet from its $N$ vertices $\{\mathbf{p}_{0},\mathbf{p}_{1},\dots,\mathbf{p}_{N-1}\}$ and construct the (N1)(N-1) linearly independent vectors that span the facet.
 $\mathbf{v}_{i}=\mathbf{p}_{i}-\mathbf{p}_{0}\quad(\text{for }i=1,\dots,N-1)$ (1)
-Next, I calculate the wedge product of these vectors.
+Next, we calculate the wedge product of these vectors.
 $\mathbf{B}=\mathbf{v}_{1}\wedge\mathbf{v}_{2}\wedge\dots\wedge\mathbf{v}_{N-1}$ (2)
-[…]
+In geometric Algebra, this (N1)(N-1)-blade $\mathbf{B}$ is a mathematical entity that represents the oriented hyperplane spanned by the vectors.
+The operation of taking the Hodge dual of this blade, $\mathbf{B}^{*}$, is equivalent to finding the orthogonal complement of that hyperplane.
+That is, for an (N1)(N-1)-dimensional hyperplane, its orthogonal complement is a 1-dimensional line, and its direction vector becomes the normal vector $\mathbf{n}$ to the hyperplane.
 $\mathbf{n}=\mathbf{B}^{*}$ (3)
@@
 Next, it is necessary to correctly determine the orientation of the facet’s normal vector.
-In conventional 3D, the correct orientation of a normal vector for a facet on the surface of a convex hull was obtained by computing the vertex winding order, which relies on the right-handed rule of the cross product.
+In the 3D case, the normal vectors of facets constituting the surface of a convex hull were obtained with correct orientation by computing the ordering of vertices, exploiting the right-handed property of the cross product of 3D vectors.
 However, the cross product is unique to 3D space.
 In higher dimensions, there are no geometric concepts analogous to the winding order that intuitively define a face’s orientation.
-To solve this problem, my implementation uses the geometric property that the normal vectors of a convex hull’s facets must always point outward.
-This orientation is determined by calculating the dot product of the facet’s normal vector and a vector pointing from a point on the facet to the centroid of the initial simplex constructed at the beginning of the algorithm.
-If the sign of the dot product is negative, the normal vector is determined to be pointing outward.
-If the sign is positive, its sign is inverted, thereby ensuring that I always obtain a normal vector with the correct outward-facing orientation.
+Therefore, in this architecture, this issue is addressed by exploiting the geometric property that facet normal vectors of a convex hull always point outward.
+The orientation is determined by computing the dot product between the facet normal vector and a vector directed from a point on the facet toward the centroid of the initial simplex constructed at the beginning of the algorithm.
+If the sign of the dot product is negative, the normal vector is judged to be outward-facing; if the sign is positive, the normal vector is flipped to obtain the correct outward orientation.
+Next, we use the facet normal vectors to determine the visible facets, which is a necessary step in the process of expanding the convex hull.
 Figure 3: The logic for determining visible facets in Quickhull.
@@
 The vector pointing from $\mathbf{P}_{\text{new}}$ to $\mathbf{P}_{\text{c}}$ is defined as $\mathbf{v}$.
-Next, I use the facet normal vectors to determine the visible facets, which is a necessary step in the process of expanding the convex hull.
 As shown in Figure 3, the facets visible from $\mathbf{P_{\text{new}}}$ can be found by examining the sign of the dot product of $\mathbf{v}$ and $\mathbf{n}$.
 The convex hull can then be expanded by connecting the new vertex $\mathbf{P_{\text{new}}}$ to the boundary of the identified set of visible facets (i.e., the set of ridges that are not shared among the visible facets).
-3.3.3 Method for Simplicial Tessellation during Convex Hull Construction
-my implementation not only computes the facets that constitute the surface of the convex hull but is also equipped with the functionality to simultaneously generate its interior volume as a set of $N$-dimensional simplices.
-This is accomplished by applying the incremental process of the Quickhull algorithm.
-In Quickhull’s incremental process, the newly added volume is represented as a set of $N$-dimensional simplices formed from the new vertex $\mathbf{p}_{\text{new}}$ and the visible facets.
-my implementation directly uses this principle, appending the corresponding simplices to a list.
-By repeating this process, I simultaneously obtain both the surface facets of the convex hull and a list of simplices that perfectly tessellates its interior volume.
-This simplicial decomposition data serves two important roles within this framework.
-Its primary application is to serve as the base data for rapidly performing a simplicial decomposition of the interior of facets clipped during Boolean operations.
-Furthermore, it is intended for future application in efficiently determining the cross-sectional structure of facets from hyperplane slicing in spaces of five or more dimensions, a topic detailed in Section 3.5.3.
+3.3.3 Simplex Filling in Convex Hull Construction
+In this architecture, Direct Quickhull is used for simplicial decomposition of the convex hull.
+In the incremental updates of Direct Quickhull, the newly added region is represented as a set of $N$-dimensional simplices generated from the new vertex $\mathbf{P}_{\text{new}}$ and the visible facets.
+Exploiting this property, the architecture appends the simplices generated at each update step to a simplex list and retains them, thereby obtaining the boundary facet set and the simplex list upon completion of convex hull construction.
+This method is used for two purposes.
+First, it is used for the simplicial decomposition of a convex polytope produced by hyperplane clipping of facets in Boolean operations.
+Second, it is intended for determining the cross-sectional structure of facets arising in hyperplane slicing in spaces of five or more dimensions; details are described in Section 3.6.3.
 3.4 N-Dimensional Boolean Operations
-3.4.1 Algorithm Overview and Design Philosophy
-While Boolean operations are an extremely powerful method for interactively modifying the shape of an $N$-dimensional mesh, their interactive implementation in high-dimensional space presents a significant challenge from a computational cost perspective.
-To address this challenge, my framework implements a practical four-stage algorithm.
-First, in a broad phase, I use spatial partitioning to narrow down potential intersection candidates.
-Next, in a narrow phase, I determine the actual geometric intersections.
-The intersected facets are then decomposed into smaller facets in a tessellation stage.
-Finally, the results are classified in an inside-outside test stage to construct the final operation result.
-In designing this algorithm, I prioritized a balance of three factors over extreme performance optimization or a full guarantee of mathematical rigor: implementation simplicity, extensibility to $N$-dimensions, and a practical level of performance that enables interactive trial and error.
-The following sections will detail the specific techniques used in each stage and the design decisions behind them.
 3.4.1 Broad Phase: Candidate Pruning via Spatial Partitioning
-In $N$-dimensional Boolean operations, a brute-force intersection test between all facets of two meshes, with $N_{\text{A}}$ and $N_{\text{B}}$ facets respectively, has a computational complexity of $O(N_{\text{A}}N_{\text{B}})$, making it a major bottleneck in interactive implementations.
-To avoid this problem, my implementation introduces a spatial partitioning grid as the broad phase to efficiently prune pairs of facets that could potentially intersect.
-I compute the axis-aligned bounding box (AABB) for each facet and register the facet with the grid cells that its AABB occupies.
-As a result, only facet pairs that share a common cell become candidates for the precise intersection tests in the subsequent narrow phase.
-Compared to hierarchical spatial partitioning methods such as BVH or Octree, the uniform grid approach adopted in my implementation offers the advantage of being simple to implement while still achieving practical pruning performance for my target use cases.
+In $N$-dimensional Boolean operations, let the numbers of facets in the two meshes be $N_{\text{A}}$ and $N_{\text{B}}$, respectively.
+A brute-force intersection test over all facet pairs then has a computational complexity of $O(N_{\text{A}}N_{\text{B}})$.
+To avoid brute-force intersection testing, this architecture adopts a spatial partitioning grid as a broad-phase stage and extracts facet pairs that are potential intersection candidates.
+We compute the axis-aligned bounding box (AABB) for each facet and register the facet with the grid cells that its AABB occupies.
+In the subsequent narrow-phase stage, intersection tests are performed on facet pairs that share the same cell.
+Hierarchical methods such as BVHs and octrees could also be considered as alternatives; however, this paper prioritizes implementation simplicity and adopts a uniform grid approach.
 Figure 4: The sequential tessellation process of a facet in a Boolean operation (analogized in 3D).
@@
 3.4.2 Narrow Phase: Facet-Facet Intersection Test
-In the narrow phase, I perform precise geometric intersection tests on the candidate pairs pruned during the broad phase.
+In the narrow phase, we perform precise geometric intersection tests on the candidate pairs pruned during the broad phase.
 The principle of this algorithm is a generalization to $N$-dimensions of the method proposed by Tomas Möller et al.
-for fast triangle-triangle intersection tests in 3D [Moller1997].
+for fast triangle-triangle intersection tests in 3D [17].
 Its essence lies in decomposing the intersection test between two facets, $\mathcal{F}_{\text{A}}$ and $\mathcal{F}_{\text{B}}$, into two symmetric sub-problems.
-First, I compute the geometric intersection region, $\mathcal{I}_{\text{B}}$, between the (N1)(N-1)-dimensional hyperplane spanned by $\mathcal{F}_{\text{A}}$ and the facet $\mathcal{F}_{\text{B}}$.
-Second, I perform the same operation with the roles swapped, computing the intersection region $\mathcal{I}_{\text{A}}$ between the hyperplane spanned by $\mathcal{I}_{\text{B}}$ and the facet $\mathcal{F}_{\text{A}}$.
+First, we compute the geometric intersection region, $\mathcal{I}_{\text{B}}$, between the (N1)(N-1)-dimensional hyperplane spanned by $\mathcal{F}_{\text{A}}$ and the facet $\mathcal{F}_{\text{B}}$.
+Second, we perform the same operation with the roles swapped, computing the intersection region $\mathcal{I}_{\text{A}}$ between the hyperplane spanned by $\mathcal{I}_{\text{B}}$ and the facet $\mathcal{F}_{\text{A}}$.
 Both of these intersection regions are (N2)(N-2)-dimensional convex polytopes.
-Finally, I determine if an intersection between the facets exists by testing whether $\mathcal{I}_{\text{A}}$ and $\mathcal{I}_{\text{B}}$ overlap within their common (N2)(N-2)-dimensional space.
+Finally, we determine if an intersection between the facets exists by testing whether $\mathcal{I}_{\text{A}}$ and $\mathcal{I}_{\text{B}}$ overlap within their common (N2)(N-2)-dimensional space.
 If an intersection is found to exist, this information is utilized in the subsequent tessellation stage.
-Specifically, I register this intersection pair in a dictionary where one facet is the key and the other is the value.
-By performing this process for all intersection pairs, I can systematically construct information about the clipping relationships, enabling us to handle complex situations where one facet intersects with multiple other facets.
-In my 4D implementation, the intersection region is a 2D convex polygon.
-Since the number of vertices is small, I adopted the Separating Axis Theorem (SAT), which is relatively straightforward to implement.
-However, in five or more dimensions, the maximum number of vertices of the (N2)(N-2)-dimensional polytope intersection region increases, making the implementation of SAT more complex.
-Therefore, for extension to higher dimensions, an intersection test approach using the GJK algorithm is considered a more practical option.
+Specifically, we register this intersection pair in a dictionary where one facet is the key and the other is the value.
+This registration is performed for all intersecting pairs, and the cutting relationships are represented as a dictionary, including cases in which a single facet has multiple intersection counterparts.
+In the 4D implementation, the Separating Axis Theorem (SAT) is adopted to determine overlap between intersection regions.
+However, in dimensions of five or higher, the number of vertices of the (N2)(N-2)-dimensional polytopes representing intersection regions can increase, which in turn increases the number of candidate separating axes enumerated by SAT and may lead to higher computational cost.
+To address this issue, for applications in five or more dimensions, an alternative candidate is intersection testing based on the GJK algorithm.
 3.4.3 Dynamic Tessellation via Direct Quickhull
 A facet that is found to intersect in the narrow phase is clipped by the (N1)(N-1)-dimensional hyperplane spanned by the opposing facet, and the clipped facet is then tessellated into multiple smaller facets.
-This entire process is illustrated in Figure 4 using an analogy in 3D.
-As the figure shows, clipping can decompose the original facet, an (N1)(N-1)-dimensional simplex, into an (N1)(N-1)-dimensional convex polytope with more than NNvertices.
-Consequently, it is necessary to efficiently re-tessellate the resulting polytope back into multiple smaller facets.
+As illustrated in Fig.
+4, a facet that is originally an (N1)(N-1)-dimensional simplex may, through cutting, be decomposed into an (N1)(N-1)-dimensional convex polytope whose number of vertices exceeds $N$.
+In such cases, the polytope after cutting must be subdivided into multiple smaller facets.
 The core of this tessellation algorithm is to classify the vertices of the facet into the positive side (P-side) and the negative side (N-side) based on their position relative to the clipping hyperplane.
 The tessellation process is executed only when the facet has vertices in both half-spaces; otherwise, no tessellation is performed.
-However, in an implementation that assumes floating-point arithmetic, this theoretical classification alone is insufficient.
-Problems can arise from numerical errors, such as an intersection that should theoretically be a single point being detected as multiple points, or the classification becoming unstable when a vertex is extremely close to the hyperplane.
-To address this challenge, my implementation introduces a two-stage vertex merging process.
-First, it unifies points that are in close proximity to each other within the computed set of intersection vertices.
-Second, it evaluates the distance between each vertex of the merged intersection region and the vertices of the original facet; if they are close, the intersection vertex is completely snapped to the coordinates of the facet vertex.
-While this operation may introduce minor deformations to the mesh geometry, it plays a crucial role in maintaining the topological consistency of the original facet within the tessellated mesh.
-After the merging process, the final tessellation is performed.
-First, I apply the simplicial tessellation feature of Direct Quickhull to the combined surface-like point cloud of the P-side vertices and the intersection region vertices to tessellate the P-side region into multiple smaller facets.
-Next, a similar operation is performed for the N-side vertices and the intersection region vertices to tessellate the N-side region.
-Through these two processes, the original facet is tessellated such that its mesh structure is consistent with the clipping plane.
-This approach demonstrates that Direct Quickhull can effectively use its capabilities even in dynamic mesh processing, thereby guaranteeing the $N$-dimensional extensibility of the entire algorithm.
+After classifying the vertices of a facet, the Direct Quickhull simplex-filling procedure is applied to a point set that combines the intersection points between the facet and the cutting hyperplane with the vertices on the positive (P) side, thereby re-subdividing the P-side region into multiple smaller facets.
+The same operation is then applied to the intersection points together with the vertices on the negative (N) side, and the N-side region is re-subdivided.
+Through these two processes, the facet is re-subdivided so as to be consistent with the cutting plane.
+In the 4D implementation, because floating-point arithmetic is assumed, numerical errors may cause issues such as detecting multiple intersection points where a single point should theoretically exist, or unstable classification when facet vertices lie extremely close to the cutting hyperplane.
+To address this issue, the following merge procedure was introduced.
+First, the intersection point set between the cutting hyperplane and the facet is computed, and duplicates are removed based on a quantization key.
+Next, among the facet vertices, points that are extremely close to the cutting hyperplane are searched.
+Let the facet vertex be $\mathbf{v}$, the normal of the cutting hyperplane be $\mathbf{n}$, an arbitrary point on the cutting hyperplane be $\mathbf{p_{\text{0}}}$, and the tolerance be $\varepsilon$.
+A point that is extremely close to the cutting hyperplane is defined as a point satisfying Eq.
+(5).
+$\lvert(\mathbf{p_{\text{0}}}-\mathbf{v})\cdot\mathbf{n}\rvert\leq\varepsilon$ (5)
+If there exists any vertex satisfying Eq.
+(5), the set consisting of the intersection point set and those vertices is treated as the intersection candidate set, and duplicates are removed again based on the quantization key.
+When an intersection point and a facet vertex are classified under the same quantization key, the intersection coordinates are replaced with the facet vertex coordinates.
+Although this operation may introduce minor deformation of the mesh, it is positioned as a process to suppress gaps and overlaps in the subdivided mesh.
 3.4.4 Inside-Outside Testing and Result Construction
@@
 This figure categorizes the general case, where a ray crosses a single facet, and representative degenerate cases that require precise intersection tests, where the ray passes exactly through a facet, edge, or vertex.
-To construct the final result of a Boolean operation—namely, the union, intersection, or difference from the set of tessellated facets, it is necessary to classify whether each facet is inside or outside the opposing object.
-For this inside-outside test, my implementation adopts the ray casting method.
-The fundamental principle of this method is based on intersection parity.
-A ray is cast from the centroid of the facet under test toward the opposing surface, and the number of intersections is counted.
-The state is then determined according to whether this number is even or odd, and the same logic applies in $N$-dimensional space.
-In my implementation, the ray is parallel to the x-axis, and its endpoint is set to a point slightly beyond the opposing mesh’s AABB.
-This ensures both computational efficiency and stability.
-To make this method operate robustly in a floating-point environment, it is necessary to handle degenerate cases where the ray passes through a facet’s vertices or edges, as shown in Figure 6.
-In an implementation that introduces a tolerance, an intersection with a vertex or edge, which should theoretically be a single point of intersection, can be detected as multiple, slightly different intersection points with the various facets that share that feature.
-This disrupts the parity of the intersection count and causes the inside-outside test to fail.
-To prevent this misclassification, my implementation employs a two-stage solution.
-First, the coordinates of the detected intersection points are discretized based on a tolerance, and multiple proximate intersection points are grouped into a single ’intersection event’ so they can be handled uniformly in subsequent processing.
+To construct the final result of a Boolean operation (union, intersection, or difference) from a set of tessellated facets, it is necessary to classify whether each facet lies inside or outside the other object.
+For inside-outside classification, this architecture adopts a ray-casting method, which is used in Boolean operations that emphasize interactivity [4].
+This method determines whether a point lies inside or outside an object based on the parity of the number of intersections between a ray cast from the point and the surface of the target object.
+In the 4D implementation, for the sake of implementation simplicity, the centroid of the facet under consideration is used as the ray origin, and the ray is cast parallel to the $x$-axis.
+The ray endpoint is set to a point slightly beyond the axis-aligned bounding box (AABB) of the opposing mesh.
+As illustrated in Fig.
+6, in degenerate cases where the ray passes through a facet vertex or edge, floating-point arithmetic may cause intersections that should theoretically be a single event at a vertex or edge to be detected as multiple, slightly different intersection points associated with multiple facets sharing that vertex or edge.
+This can lead to misclassification in inside-outside tests.
+To address this issue, we quantize the detected intersection points using a tolerance and group points with the same quantization key into a single intersection event.
 Next, the nature of this grouped intersection event is determined.
@@
 The left diagram shows a piercing case, and the right diagram shows a tangent case.
-To do this, I evaluate the dot product of the ray’s direction vector and the normal vector of every facet included in the event.
+To do this, we evaluate the dot product of the ray’s direction vector and the normal vector of every facet included in the event.
 If the signs of the dot products are biased to one side (either all positive or all negative), the event is considered to be piercing the surface, and the intersection count is incremented by one.
@@
 Based on this inside-outside test executed through this series of processes, the tessellated facets are classified according to the type of Boolean operation, and the final mesh is constructed.
-However, in cases where the ray is coplanar with a facet, such as the ”Co-planar Facet” case shown in Figure 5, the dot product of the ray and the normal vector is theoretically zero, causing the sign-based test to fail robustly.
-Therefore, there is no guarantee that intersections can always be correctly determined under these conditions.
-A practical approach would be to avoid such degenerate geometric configurations by using existing techniques, such as ray perturbation.
-3.5 Visualization of N-Dimensional Objects
-3.5.1 Cross-Sectioning as the Basic Strategy
-This component is responsible for generating and rendering 3D cross-sections based on $N$-dimensional mesh data.
-The basic strategy to achieve this transformation is based on the concept of sequentially reducing the dimension.
-This is a process of recursively clipping an $N$-dimensional object with an (N1)(N-1)-dimensional hyperplane, and then clipping the resulting cross-section with an (N2)(N-2)-dimensional hyperplane.
-In this paper, I refer to this method of sequentially reducing dimensions as Hierarchical Cross-Sectioning.
-With this approach, the complex problem of computing cross-sections in arbitrary dimensions can always be reduced to the simpler problem of the intersection between an $N$-dimensional facet and an (N1)(N-1)-dimensional hyperplane.
-In the following sections, as a concrete application of this Hierarchical Cross-Sectioning for $N=4$, I will detail the specific algorithm for generating 3D cross-sections from 4D objects and discuss its implementation challenges.
-3.5.2 Facet Clipping Algorithm in the 4D Implementation
-To generate a 3D cross-section of an object in 4D space, my implementation computes the geometric intersection between each facet of the 4D object—a tetrahedron—and an axis-parallel hyperplane $w=w_{\text{slice}}$ (where $w=w_{\text{slice}}$ is an arbitrary constant).
-This procedure is executed independently for every facet of the input mesh according to the following steps.
-First, I determine on which side of the clipping hyperplane each of the facet’s four vertices is located.
-If all vertices are on the same side, the facet does not intersect the hyperplane and skip intersection computation.
-If the vertices straddle the hyperplane, the intersection point of each constituent edge of the facet with the hyperplane is calculated via linear interpolation.
-Let the two vertices defining an edge be
-$\mathbf{p_{0}}=({x_{0}},{y_{0}},{z_{0}},{w_{0}})\quad\mathbf{p_{1}}=({x_{1}},{y_{1}},{z_{1}},{w_{1}})$ (5)
-The interpolation parameter ttfor determining the position of the intersection point is then calculated from the $w$-coordinates of both vertices and the $w$-coordinate of the clipping plane, $w_{\text{slice}}$, as shown in Equation (6).
-$t=\frac{w_{\text{slice}}-w_{0}}{w_{1}-w_{0}}$ (6)
-Next, using this parameter $t$, the coordinates of the intersection point $\mathbf{p}_{\text{isect}}$ in 4D space are calculated according to Equation (7).
-$\mathbf{p}_{\text{isect}}=\mathbf{p}_{1}+t(\mathbf{p}_{1}-\mathbf{p}_{0})$ (7)
-The 3D coordinates, obtained by extracting the $x,y,z$ components of the resulting 4D coordinate vector $\mathbf{p}_{\text{isect}}$, directly become the coordinates of a vertex that constitutes the cross-section.
-By performing this operation for all edges that intersect the hyperplane, I obtain the set of vertices of the cross-section, which are then connected to form a polygon.
-The shape of the cross-section generated by the intersection of a single 4D facet and a 3D hyperplane is, within 3D space, either a triangle with three vertices or a convex polygon with four vertices.
-A 4-vertex convex polygon can be divided into two triangles to obtain a set of triangular polygons for rendering.
-3.5.3 Simplicial Decomposition of the Cross-Section
-Figure 7: Dimensional dependence of the number of cross-section vertices generated during hyperplane clipping.
-When clipping a facet in 3D, the number of vertices is always two.
-However, when clipping a 3D facet in 4D, the number varies between three and four depending on the relative position of the hyperplane, and is thus not uniquely determined.
-Regarding the color coding in this figure, the blue plane represents the clipping hyperplane, and the red area indicates the portion of it that exists on the dimension being clipped.
-The darker shades are highlights for visual clarity.
-The cross-section generated by the above clipping algorithm is generally an (N2)(N-2)-dimensional convex polytope.
-For subsequent processing, this must be decomposed into a set of (N2)(N-2)-dimensional simplices.
-In the 4D implementation, the cross-section is a 2D convex polygon (either a triangle or a quadrilateral) with a maximum of four vertices, as shown in Figure 7.
-Although the order of the vertex list is undefined, the polygon can be safely tessellate into two triangles by sorting the vertices by angle around the polygon’s centroid and then systematically partitioning it.
-my implementation efficiently performs this sort using only vector operations, avoiding computationally expensive trigonometric functions.
-However, this angle-sorting method cannot be generalized to five or more dimensions $(N\geq 5)$.
-The maximum number of vertices of an (N2)(N-2)-dimensional cross-section, obtained by clipping an $N$-dimensional facet (which has NNvertices) with a hyperplane, is given by the well-known result in combinatorics shown in Equation (8).
-$\left\lfloor\frac{N}{2}\right\rfloor\cdot\left\lceil\frac{N}{2}\right\rceil$ (8)
-For example, in 5D, the cross-section can be a 3D convex polytope with a maximum of $\lfloor 5/2\rfloor\cdot\lceil 5/2\rceil=6$ vertices.
-The simplicial tessellation of such a complex polytope is a non-trivial computational problem.
-As a dimension-independent and general solution to this high-dimensional extensibility challenge, I can use the simplicial tessellation feature of Direct Quickhull, which was detailed in Section 3.3.3.
-By providing the set of vertices that constitute the cross-section as input to Direct Quickhull, it is possible to efficiently tessellate the interior of the cross-sectional polytope itself into a set of (N1)(N-1)-dimensional facets.
-As mentioned in Section 3.3.1, my implementation, which omits the interior point search, is well-suited for processing this type of surface-like point cloud.
-This ensures that the visualization pipeline of my framework is, in principle, extensible to higher dimensions.
-3.5.4 Robust Method for Determining Cross-Section Normals
-When generating a 3D cross-section for rendering from a 4D object, it is necessary to determine the winding order of the vertices that form the 3D facet.
-The clipping algorithm from the previous section produces a list of vertices in an undefined order.
-If this unordered vertex list is used as is, a serious problem arises: many rendering pipelines determine the front and back of a face based on the input order, which would result in inconsistent face orientations and prevent the correct rendering of the shape.
-To solve this problem, my method utilizes the normal vector of the original 4D facet as a reference for the correct orientation.
-Specifically, I first project the original 4D normal vector into 3D space by extracting its $x,y,z$ components, and define this as the true normal vector that the 3D facets should follow.
-Next, I compute a temporary normal vector from the generated vertex set using the cross product and evaluate its dot product with the true normal vector.
-If the sign of this dot product is negative, the winding order of the vertices is reversed.
-This ensures that all triangles constituting the cross-sectional mesh have a consistent vertex winding order that can be correctly interpreted by a general-purpose renderer.
-3.6 Pose Representation and Control Scheme for Interactive Exploration
+However, in cases where the ray lies on the same plane as a facet, such as the “co-planar facet” case illustrated in Fig.
+5, the dot product between the ray direction and the facet normal vector is theoretically zero.
+As a result, misclassification may occur in sign-based decision methods.
+Therefore, under such conditions, there is no guarantee that intersections can always be classified correctly.
+To avoid these degenerate geometric configurations, techniques such as ray perturbation can be considered as candidate approaches.
+3.5 Coordinate Transformations for High-Dimensional Exploration
 3.5.1 Pose Representation and Management with Geometric Algebra Rotors
-In representing the pose of an $N$-dimensional object, conventional 3D methods like Euler angles or quaternions cannot be directly applied due to issues such as gimbal lock and difficulties in extending them to higher dimensions.
-To solve this challenge, my framework adheres to the mathematical foundation of Geometric Algebra, as mentioned in Section 3.2, and represents the object’s pose uniformly using a Rotor.
+In this architecture, the orientation of an $N$-dimensional object is represented using rotors, in accordance with geometric algebra, which serves as the mathematical foundation described in Section 3.1.
 A rotor is a rotation operator defined by a specific plane of rotation and a rotation angle, which allows the composition of rotations and their application to vectors to be handled as an algebraic product, regardless of the dimension.
@@
 Here, $\tilde{R}$ is the reverse of the rotor $R$. If $R$ is normalized, $\tilde{R}$ is equivalent to $R^{-1}$.
-In my implementation, instead of having the user manipulate the rotor directly, I allow them to specify the amount of rotation for each of the six fundamental rotation planes in 4D space $(xy,xz,xw,yz,yw,zw)$ as individual parameters, similar to Euler angles.
-These parameter values are then composed in a predefined order to internally construct the final rotation operator, the rotor.
-This design, which separates the intuitive parameters provided to the user from the mathematically robust operator used internally, completely avoids gimbal lock and is easily extensible to higher dimensions.
-Furthermore, my implementation fixes the clipping hyperplane to be axis-parallel $(w=\text{const})$ and transforms the object’s pose instead.
-From the user’s perspective, this approach is visually equivalent to clipping a stationary object with an arbitrarily oriented hyperplane.
-However, it replaces the computationally expensive problem of arbitrary hyperplane clipping with a combination of an efficient pose transformation and an axis-parallel clip, enabling interactive observation of the cross-sections.
-3.6.2 Camera-Based Coordinate Transformation
+In the 4D implementation, users do not manipulate rotors directly; instead, rotation amounts are specified as individual parameters, analogous to Euler angles, for each of the six fundamental rotation planes in 4D space $(xy,xz,xw,yz,yw,zw)$.
+These parameter values are composed in a prescribed order to internally construct a rotor as the final rotation operator, which is then used for rotation computation.
+3.5.2 View Transformation
 This section defines the view transformation in $N$-dimensions, which transforms vertices from the world coordinate system to the camera coordinate system based on the camera’s state (position and pose).
@@
 In this setup, the camera’s position $\mathbf{C}$ corresponds to the origin of the camera coordinate system.
-To intuitively understand the geometric essence of this transformation, I consider the 2D space shown in Figure 8 as an example.
-I denote the position of a target vertex $\mathbf{P}$ in the world coordinate system as $\mathbf{P}^{(\mathrm{W})}$ and the position of the camera $\mathbf{C}$ as $\mathbf{C}^{(\mathrm{W})}$.
+To intuitively understand the geometric essence of this transformation, we consider the 2D space shown in Figure 7 as an example.
+We denote the position of a target vertex $\mathbf{P}$ in the world coordinate system as $\mathbf{P}^{(\mathrm{W})}$ and the position of the camera $\mathbf{C}$ as $\mathbf{C}^{(\mathrm{W})}$.
 The coordinate transformation takes $\mathbf{P}^{(\mathrm{W})}$, $\mathbf{C}^{(\mathrm{W})}$, and the rotor $R$ representing the pose as input.
-The transformation process first derives the basis vectors that span the camera coordinate system, $\mathbf{e}_{\text{right}}^{(\mathrm{W})}$ and $\mathbf{e}_{\text{fwd}}^{(\mathrm{W})}$, from the rotor $R$ representing the camera’s pose, as shown in Equation (11).
+In the transformation process, the rotor $R$ representing the camera orientation and the basis vectors of the world coordinate system are first used to rotate the basis vectors according to Eq.
+(8), thereby deriving the basis vectors $\mathbf{e}_{\text{right}}^{(\mathrm{W})}$ and $\mathbf{e}_{\text{fwd}}^{(\mathrm{W})}$ that span the camera coordinate system.
 Next, the relative position vector $\mathbf{v}^{(\mathrm{W})}$ from the camera’s position to the object’s vertex is calculated as shown in Equation (9).
@@
 By using homogeneous coordinates, these two operations can be combined into a single $(N+1)\times(N+1)$ matrix, analogous to the view matrix commonly used in computer graphics.
-The description using dot products in this paper is specifically intended to directly illustrate the geometric meaning of the rotation operation.
-3.6.3 Design of the Control Scheme for Interactive Exploration
-To address the challenge of mapping the six independent rotational degrees of freedom required to control the camera’s orientation in 4D space onto a standard mouse and keyboard, I adopted two design principles: orthogonality of controls and leveraging existing control conventions (specifically, those of First-Person Shooters, or FPS).
-The 3D translations (WASD keys) and pitch rotation (mouse vertical movement) follow standard conventions.
-The primary challenge, which is the assignment of a rotation to the horizontal mouse movement ($x$-axis), is resolved by a modal input switch using a modifier key (Left Control).
-In its normal mode, horizontal mouse movement controls the primary yaw rotation ($xz$-plane) for 3D exploration.
-While the modifier key is held down, this input is switched to control the rotation in the $xw$-plane, which serves a similar role in 4D exploration.
-This design ensures that a single input always corresponds orthogonally to a single operation, resulting in a highly predictable control scheme that allows the user to operate while being clearly conscious of switching between 3D and 4D exploration modes.
-3.6.4 Extension of the Manipulation System to N Dimensions
-This design principle can be generalized to $N$-dimensions.
-I apply a systematic approach to address the rotational degrees of freedom in an $N$-dimensional space, which increase quadratically as N(N1)/2N(N-1)/2.
-First, translations and rotations within the 3D cross-section that the user directly observes are kept on the standard control scheme (WASD/mouse), regardless of the total number of dimensions.
-This provides the user with a consistently stable operational base.
-Next, for the remaining (N(N1)/2)3(N(N-1)/2)-3 rotations outside the observed space, I consider an approach that generalizes the design of my 4D implementation.
-Specifically, horizontal mouse movement is consistently assigned to rotations that occur outside the observed space; the user selects the target high-dimensional axis $(w,u,\ldots)$ via a combination of modifier keys, which in turn manipulates the corresponding rotation plane $(xw,xu,\ldots)$.
-Because this approach dedicates the primary rotational input to high-dimensional exploration while allowing the control scheme to be systematically extended, it is expected to provide a scalable foundation for exploring $N$-dimensional space while preserving control orthogonality.
-3.7 Demonstration of Extensibility
-The core of this framework’s extensibility lies in the design principle of separating topology and geometry.
-This section demonstrates the effectiveness of this design through a case study in which a 4D non-rigid body physics simulation is integrated.
-A practical challenge in this integration is the mismatch of data formats.
-For computational efficiency, the physics simulation requires unique, non-duplicated vertex data.
-The rendering pipeline, on the other hand, requires duplicated vertex data to accommodate correct normals and UVs.
-The separated architecture of my framework resolves this discrepancy through index mapping.
-During initialization, the physics engine extracts the set of unique vertices from the rendering mesh’s geometry and builds a map, just once, that maintains the correspondence between the indices of the unique vertices and the indices of the original, duplicated vertices.
-During the simulation, computations are performed only on the set of unique vertices.
-Then, in each frame, this map is used to efficiently update the entire vertex array for rendering.
-This pattern—maintaining an internally optimized data format and converting it to the required format only during communication is a general solution for integrating any external component that requires a different data format, and is not limited to this case study.
-This demonstrates that the design principle of separating topology and geometry is what enables the high modularity and extensibility of my framework.
-4 Plex: An N-Dimensional Mesh Data Exchange Ecosystem
+The dot-product-based description in this paper is intended to make the geometric meaning of the rotation operation explicit.
+3.6 Visualization of N-Dimensional Meshes
+3.6.1 Cross-Sectioning as the Basic Strategy
+This component generates and renders 3D cross-sections from $N$-dimensional mesh data.
+The basic strategy for realizing this process is based on the idea of sequentially reducing dimensionality.
+This is a process of recursively clipping an $N$-dimensional object with an (N1)(N-1)-dimensional hyperplane, and then clipping the resulting cross-section with an (N2)(N-2)-dimensional hyperplane.
+In this paper, we refer to this method of sequentially reducing dimensions as Hierarchical Cross-Sectioning.
+With this approach, complex cross-section computation problems in arbitrary dimensions are always reduced to a simpler problem: the intersection between an $N$-dimensional facet and an (N1)(N-1)-dimensional hyperplane.
+However, because the information obtained through sequential hyperplane slicing is limited to cross-sections, it involves information loss, and the original shape cannot be uniquely reconstructed from the cross-sections alone.
+A quantitative evaluation of such perceptual limitations is beyond the scope of this paper.
+In the following sections, as a concrete application of this Hierarchical Cross-Sectioning for $N=4$, we will detail the specific algorithm for generating 3D cross-sections from 4D objects and discuss its implementation challenges.
+3.6.2 Facet Clipping Algorithm in the 4D Implementation
+To generate 3D cross-sections of objects embedded in 4D space, the present implementation computes the geometric intersections between each facet of the 4D mesh, represented as tetrahedra, and an axis-aligned hyperplane $w=w_{\text{slice}}$, where $w_{\text{slice}}$ is an arbitrary constant.
+This computation is performed for all facets of the input mesh.
+Cross-section generation based on this type of slicing plane has also been adopted in existing studies on interactive 4D visualization [5].
+In this implementation, no mechanism is provided to arbitrarily specify the orientation of the cutting hyperplane; cross-sections are always generated using the axis-aligned plane $w=w_{\text{slice}}$.
+When observing the cross-sections from arbitrary directions, changes in the viewing direction are handled as view transformations.
+First, we determine on which side of the clipping hyperplane each of the facet’s four vertices is located.
+If all vertices are on the same side, the facet does not intersect the hyperplane, and we skip the intersection computation.
+If the vertices straddle the hyperplane, the intersection point of each constituent edge of the facet with the hyperplane is calculated via linear interpolation.
+Let the two vertices defining an edge be
+$\mathbf{p_{0}}=({x_{0}},{y_{0}},{z_{0}},{w_{0}})\quad\mathbf{p_{1}}=({x_{1}},{y_{1}},{z_{1}},{w_{1}})$ (11)
+The interpolation parameter $t$ for determining the position of the intersection point is then calculated from the $w$-coordinates of both vertices and the $w$-coordinate of the clipping plane, $w_{\text{slice}}$, as shown in Equation (12).
+$t=\frac{w_{\text{slice}}-w_{0}}{w_{1}-w_{0}}$ (12)
+Next, using this parameter $t$, the coordinates of the intersection point $\mathbf{p}_{\text{isect}}$ in 4D space are calculated according to Equation (13).
+$\mathbf{p}_{\text{isect}}=\mathbf{p}_{0}+t(\mathbf{p}_{1}-\mathbf{p}_{0})$ (13)
+The 3D coordinates obtained by extracting the $x,y,z$ components of the resulting 4D coordinate $\mathbf{p}_{\text{isect}}$ are directly used as the vertex coordinates constituting the cross-section.
+By performing this operation for all edges of facets that intersect the hyperplane, the vertex set of the cross-section is obtained, and polygons are constructed by connecting these vertices.
+The shape of the cross-section generated by the intersection of a single 4D facet and a 3D hyperplane is, within 3D space, either a triangle with three vertices or a convex polygon with four vertices.
+For a convex polygon with four vertices, it is divided into two triangles to obtain a set of triangular polygons for rendering.
+3.6.3 Simplicial Decomposition of the Cross-Section
+Figure 8: The dimensional dependence of the number of cross-section vertices generated by hyperplane cutting is as follows.
+In three dimensions, facet cutting produces two vertices, whereas in four dimensions the number of vertices becomes three or four depending on the hyperplane’s relative position.
+Regarding the color coding in this figure, the blue plane represents the clipping hyperplane, and the red area indicates the portion of it that exists on the dimension being clipped.
+The darker shades are highlights for visual clarity.
+The cross-sections generated by the aforementioned cutting algorithm form an (N2)(N-2)-dimensional convex polytope in $N$ dimensions.
+For subsequent processing, this must be decomposed into a set of (N2)(N-2)-dimensional simplices.
+The maximum number of vertices of an (N2)(N-2)-dimensional cross-section obtained by cutting an $N$-dimensional facet (with $N$ vertices) by a hyperplane is given by  (14).
+$\left\lfloor\frac{N}{2}\right\rfloor\cdot\left\lceil\frac{N}{2}\right\rceil$ (14)
+For example, in 5D, the cross-section of a facet can be a 3D convex polytope with a maximum of $\lfloor 5/2\rfloor\cdot\lceil 5/2\rceil=6$ vertices.
+For simplicial decomposition of such polytopes, the simplex-filling function of Direct Quickhull described in Section 3.3.3 is used.
+By providing the vertex set constituting the cross-section as input to Direct Quickhull, the cross-section is subdivided into a set of (N2)(N-2)-dimensional simplices.
+In the 4D implementation, as illustrated in Fig.
+8, the cross-section is a two-dimensional convex polygon with at most four vertices.
+Therefore, the vertex order is arranged by angle sorting using the cross product, and a four-vertex two-dimensional convex polygon is subdivided into two triangles.
+3.6.4 Determination of Cross-Section Normals and Vertex Winding Order
+In the 4D implementation, when generating 3D cross-sections to be rendered from a 4D object, it is necessary to determine the vertex winding order of the 3D facets.
+The cutting algorithm described in the previous section produces an unordered set of vertices.
+If this unordered vertex set is used as is, many rendering pipelines determine front-facing and back-facing polygons based on the input order, which can result in inconsistent face orientations and, consequently, incorrect rendering of the shape.
+To address this issue, the present implementation uses the normal vector of the original 4D facet as a reference for the correct orientation.
+First, the $x,y,z$ components of the 4D facet normal vector are extracted to project the normal vector into 3D space, and this projected vector is defined as the reference normal vector for determining the winding order of the 3D facet.
+Next, the dot product between this reference normal vector and a provisional normal vector computed by the cross product from the vertex set of the generated 3D facet is evaluated.
+If the sign of the dot product is negative, the vertex winding order is reversed, thereby ensuring consistent winding of the triangles constituting the cross-section mesh.
+3.7 Interaction Scheme for Interactive Exploration
+3.7.1 Interaction Scheme Design for the 4D Implementation
+In the 4D implementation, among the six rotational degrees of freedom, the operations handled by the mouse and keyboard are assigned according to the following policy.
+First, to keep the mapping between inputs and actions as orthogonal as possible, we adopt FPS navigation as the basis, separating translation to the keyboard (WASD) and viewpoint rotation to the mouse.
+Navigation centered on continuous first-person movement has been evaluated in comparative studies of interaction techniques for exploration tasks [21, 16].
+Second, to handle high-degree-of-freedom rotations with limited 2D input, the assignment of horizontal mouse movement ($x$-axis) is switched using a modifier key (LeftControl).
+In the default state, it is assigned to yaw rotation (the $xz$ plane), and while the modifier key is held down, it is switched to rotation in the $xw$ plane.
+Such a “spring-loaded mode” (quasimode), which is active only while the key is held, has been discussed as a design that can reduce mode errors compared to persistent modes [11].
+3.7.2 Application of Interaction Design Policies to N Dimensions
+In this section, we describe an interaction scheme obtained by applying the assignment policy used in the aforementioned 4D implementation to $N$ dimensions.
+Since the rotational degrees of freedom in an $N$-dimensional space increase to N(N1)/2N(N-1)/2, translation and rotation within the observed 3D cross-section are designed to use dimension-independent operations (WASD/mouse).
+In contrast, for rotations outside the observed space, following the same principle as in the 4D implementation, additional degrees of freedom are introduced incrementally while preserving input orthogonality.
+Specifically, horizontal mouse movement is assigned to rotations outside the observed space, higher-dimensional axes $(w,u,\ldots)$ are selected via combinations of modifier keys, and the corresponding rotation planes $(xw,xu,\ldots)$ are designated as the manipulation targets.
+This type of switching is consistent with a design policy in which constraints are used to control the degrees of freedom handled by the user [20].
+3.8 Design Based on Topological and Geometric Separation
+In physical simulations and related applications, when handling geometry that changes over time while sharing the same topology, a unique vertex array constituting the mesh is typically assumed as input.
+In contrast, visualization and rendering pipelines treat facets as the minimal unit; therefore, when a collection of facets is represented using vertex sets, duplicate indices arise, and such pipelines assume vertex arrays that include these duplicates as input.
+As a result, the data formats required by update systems, which demand unique vertex arrays, and by visualization and rendering systems, which assume vertex arrays containing duplicated indices, do not coincide.
+Under the assumptions of the update and visualization/rendering systems considered in this study, it is therefore not possible to integrate both within a single architecture without separating topology and geometry.
+To address this issue, the proposed architecture adopts a design in which, at the initialization stage, a correspondence table (index mapping) between a unique vertex array and a vertex array containing duplicated indices is maintained, derived from the mesh data.
+This design allows update systems that take a unique vertex array as input and visualization/rendering systems that assume vertex arrays with duplicated indices to be handled simultaneously within the same architecture.
+Update processing is completed solely on the unique vertex array, while reflection to visualization and rendering is separated as a transfer process based on the correspondence table.
+In the 4D implementation, based on this design, a non-rigid body simulation based on XPBD is integrated, and it is confirmed that mesh vertex coordinates are updated over time and reflected in cross-section generation and rendering.
+However, evaluation of the validity or accuracy of the physical model is not included in the scope of this study.
+4 Plex: An N-Dimensional Mesh Data Exchange Format
 Table 1: Overview of standard chunks in the .plex format
@@
 ’FACE’ Required Stores the vertex indices (topology) that constitute the facets.
-’NORM’/’NORMD’ Required Stores the normal vectors that define the orientation of the facets in single or double precision.
-’CENT’/’CENTD’ Optional Stores the cached centroid coordinates of facets to reduce computational load.
+’NORM’/’NRMD’ Required Stores the normal vectors that define the orientation of the facets in single or double precision.
+’CENT’/’CNTD’ Optional Stores the cached centroid coordinates of facets to reduce computational load.
 Table 2: Internal data layout of the ’META’ chunk
@@
 0x28 Variable char[] Software Name: UTF-8 encoded string.
-The $N$-dimensional simulation environments targeted by this research require, as their foundation, an extensible data exchange ecosystem for uniformly handling high-dimensional mesh data.
-However, to date, no widely accepted standard format for exchanging $N$-dimensional mesh data has existed, leading to a fragmented situation where formats vary among researchers and projects.
-This state of affairs has been a barrier to the smooth sharing of data between different tools and research groups.
-To address this challenge, this study defines a systematic data exchange ecosystem for $N$-dimensional meshes, named ”Plex” (with the .plex extension), which is a file format featuring a chunk-based architecture.
+4.1 Background and Challenges
+In the $N$-dimensional simulation environments targeted by this study, there are cases where a means for exchanging high-dimensional mesh data between projects is required.
+However, for $N$-dimensional meshes, there are only a limited number of common exchange formats that are widely used across the research community.
+In practice, project-specific formats or extensions of existing formats are often employed.
+This situation can lead to additional effort when sharing data across different implementations or research groups, including data conversion, validation, and interpretation of specifications.
+To address this issue, this study defines a chunk-based file format, “Plex” (.plex), for the purpose of exchanging $N$-dimensional mesh data.
+4.2 Relation to Existing Formats and Positioning of Plex
+HDF5 and netCDF are formats designed to store multidimensional array data in a self-describing manner and can handle $N$-dimensional datasets [9, 19].
+However, because their primary purpose is array storage, they do not directly specify the topology and geometry of $N$-dimensional meshes as part of their specifications.
+The VTK file formats are organized as a family of formats for describing datasets that include points, cell elements, and their associated attributes [13].
+XDMF adopts a model for representing data such as meshes based on XML-based metadata descriptions and references to external data [25].
+In contrast, this study focuses on defining, as a single specification, the minimal elements commonly handled by mesh data parameterized by the dimension $N$.
+These elements are topology and geometry.
+This focus does not align with the design objectives of these formats.
+Based on the above considerations, this study does not aim to replace existing formats; instead, Plex is positioned as a format that explicitly describes the minimal elements (topology and geometry) commonly represented in $N$-dimensional mesh data.
 4.3 Dual-Format Strategy
-Research workflows that handle $N$-dimensional data present two distinct requirements: the need to inspect and verify data content in the early stages, and the need for high processing performance in large-scale computations.
-To meet these requirements, Plex adopts a dual-format strategy, following the design philosophy proven by the glTF format in the 3D domain.
-The first format is a schema based on the standard JSON technology, which prioritizes data readability.
-This allows for the easy inspection and debugging of data content using a rich ecosystem of existing tools.
-The second is a chunk-based binary format, .plex, which prioritizes execution performance.
-This format is intended for fast loading by applications and efficient data storage.
-The reference implementation provided in this study includes a one-way converter from the .plex format to the JSON format.
-This enables researchers to typically work with the high-performance binary format while exporting to a human-readable format as needed, thereby achieving a balance between execution efficiency and data transparency.
+In research workflows that handle $N$-dimensional data, two different requirements may coexist: inspection and validation of data contents at early stages, and processing performance for large-scale computations.
+To address these requirements, Plex adopts a design that combines a human-readable JSON format and a chunk-based binary format .plex.
+The human-readable format is used for content inspection and debugging, while the binary format is used to improve efficiency in loading and saving.
+The present implementation includes a one-way converter from .plex to the JSON format.
+This enables a workflow in which data are normally stored in the binary format and exported to the human-readable format only when necessary.
+In this paper, this conversion is presented as a minimal means for content inspection and third-party verification of Plex data, while providing converters to VTK-family formats or HDF5 is considered out of scope and positioned as future work.
 4.4 .plex Binary Format Specification
-The .plex format is designed with a chunk-based architecture, following the proven extensibility model of standard formats such as PNG.
-A file consists of a global header followed by a sequence of at least four self-descriptive chunks as shown in Table 1; the internal structure of the META chunk is detailed in Table 2.
-Each chunk contains a 4-character chunk type, the data length, the data payload, and a CRC32 checksum to verify data integrity.
-The data payload structure for the primary chunks, excluding the META chunk, is defined as follows.
-The payload of the FACE chunk, which defines the topology, first specifies the total number of facets as a 64-bit unsigned integer, followed by a list of vertex indices (an array of 32-bit unsigned integers) that constitute each facet.
-In contrast, the payload for the VERT or VERD chunks, which store vertex coordinates, contains no metadata such as counts; for a mesh with $V$ vertices in $N$ dimensions, it consists solely of a contiguous array of $V\times N$ numerical values.
-The respective data types are 32-bit single-precision and 64-bit double-precision floating-point numbers.
-The payload layout for the NORM/NRMD chunks (storing facet normals) and the optional CENT/CNTD chunks (storing centroids) also conforms to this VERT/VERD chunk format.
-A parser compliant with the specification is guaranteed to safely skip any unknown chunk types it cannot interpret.
-This ensures that backward compatibility with existing software is maintained even if the community adds new data in the future, such as textures or animations, as custom chunks.
-Furthermore, to encourage community-driven extensions, a naming convention is established that reserves identifiers containing one or more lowercase letters for use as custom chunks.
-This standardization will facilitate data sharing among researchers and serve as a technical foundation for improving the reproducibility of published results.
-Moreover, the forward compatibility guaranteed by this specification allows for future extensions, providing a sustainable infrastructure that transforms $N$-dimensional mesh data from something previously confined to individual environments into a resource that can be shared and reused by the entire community.
+The .plex format is defined as a chunk-based format consisting of a sequence of self-describing chunks.
+A file comprises a global header and at least four self-describing chunk sequences shown in Table 1, and the internal structure of the META chunk is shown in Table 2.
+Each chunk contains a four-character chunk type, a data length, a data payload, and a CRC32 checksum for verifying data integrity.
+The data payloads of the primary chunks, except for the META chunk, are defined as follows.
+In the FACE chunk, which defines topology, the total number of facets is recorded as a 64-bit unsigned integer, followed by the vertex index lists (arrays of 32-bit unsigned integers) that constitute each facet.
+The VERT or VERD chunks that store vertex coordinates contain no metadata such as counts; for an $N$-dimensional mesh with $V$ vertices, they consist solely of a contiguous array of $V\times N$ numerical values.
+The data types are 32-bit single precision and 64-bit double precision, respectively.
+The data layouts of the NORM/NRMD chunks and optional CENT/CNTD chunks likewise conform to the formats of the VERT/VERD chunks.
+From a compatibility perspective, parsers compliant with the specification are required to be able to skip unknown chunk types, allowing loading to continue as long as the existing required chunks can be interpreted.
+In addition, assuming future extensions such as additional attributes or auxiliary information to be introduced as chunks, a naming convention is defined in which identifiers containing at least one lowercase letter are reserved for custom chunks.
+Here, “compatibility” refers to forward compatibility within the Plex specification, whereby loading can proceed using only known chunks while ignoring unknown ones, and does not imply binary compatibility with VTK-family formats or HDF5.
 Figure 9: The set of base objects used in the experiments.
 The top row shows the 3D objects, and the bottom row shows the 4D objects.
-The number to the right of each object’s name is the number of facets that constitute its mesh.
-The 4D objects are visualized using the hyperplane slicing feature of my framework, with their 3D cross-sections at three representative $w$-coordinates (w=0.8,0,0.8)(w=-0.8,0,0.8) displayed side-by-side.
+The numerical value shown to the right of each name indicates the mesh size (number of facets).
+4D objects are displayed as 3D cross-sections obtained using the hyperplane slicing function described in this paper.
+As an example, cross-sections at three $w$ coordinates (w=0.8,0,0.8)(w=-0.8,0,0.8) are arranged from left to right.
 5 Experiments and Evaluation
 5.1 Experimental Setup
-The performance evaluation in this section is not intended to be a speed comparison with existing offline methods.
-Instead, its purpose is to evaluate the performance scalability of the proposed CPU-based framework with respect to the scale of the processed mesh (e.g., the number of facets).
+The performance evaluation in this chapter, for the 4D implementation, does not aim to compare execution speed with existing offline methods.
+Instead, it aims to evaluate the scalability of performance with respect to the size of the target meshes (e.g., the number of facets) within the proposed CPU-based architecture.
 All implementation and performance evaluations in this study were conducted in the following environment.
@@
 Table 4: Performance evaluation results for Boolean operations.
-I performed union, intersection, and difference operations in 3D and 4D spaces using combinations of the base objects shown in Figure 9.
-For each combination, the table shows the processing time and the size (number of facets) of the resulting mesh.
-The processing time is the average of three trials.
+For combinations of the basic objects shown in Fig.
+9, union, intersection, and difference operations were performed in three- and 4D spaces.
+Each entry reports the processing time averaged over three trials ($\mathrm{ms}$) and the size of the generated mesh (number of facets).
 Input Union Intersection Difference
@@
 Table 5: Interactive performance evaluation.
-This table shows the rendering performance in scenes containing 1 to 20 4D objects (hypercubes with 47 facets each), while the user continuously performed movement and view rotation operations in the xzxz and xwxw planes for 30 seconds.
-Each value is the average of three 30 second measurement trials.
+Rendering performance is shown for a scene in which 1–20 4D hypercubes, each with 47 facets, are placed, while continuous interaction (translation and viewpoint rotation in the xzxz and xwxw planes) is performed for 30 seconds.
+As performance metrics, the average FPS, minimum FPS, and the 99th percentile frametime ($\mathrm{ms}$) are reported.
+All values are averages over three 30-second measurement runs.
 Objects (Total Facets) Average FPS Minimum FPS 99th Percentile Frametime ($\mathrm{ms}$)
@@
 20(940) 21.2 13.1 58.1
-5.2.1 Performance of the Direct Quickhull Implementation
-As shown in Table 3, for small-scale point clouds with 50 or fewer vertices which are the intended primary use case in my framework the processing time is under 1.7 $\mathrm{ms}$ in either dimension, confirming that its performance is sufficient for this purpose.
-5.2.2 Performance of N-Dimensional Boolean Operations
-Next, I evaluated the performance of the $N$-dimensional Boolean operations, which are central to interactive mesh editing.
-I executed union, intersection, and difference operations between various objects in 3D and 4D spaces and measured their computation times.
-The results are presented in Table 4.
-In 3D space, the operations were generally completed in several tens of milliseconds, depending on the combination of objects and the type of operation.
-In 4D space, however, the computational load increased significantly.
-Even a simple combination of two Cubes required approximately 300 $\mathrm{ms}$, and in some cases, operations between two complex Random objects took over two seconds to complete.
-These results indicate that high-dimensional Boolean operations can be executed on a time scale of a few seconds.
-This processing time is considered to be within an acceptable range for an interactive workflow, where the user waits for the result of an operation before considering the next action.
-This demonstrates the feasibility of incorporating these operations, which were previously dominated by offline batch processing, into interactive applications.
-5.3 Interactive Performance Evaluation
-This section quantitatively evaluates the performance characteristics of the proposed framework.
-I placed multiple simple 4D objects, each with 47 facets, and measured various performance metrics as the total number of facets in the scene was increased.
-The experimental results are shown in Table 5.
-These results indicate that the performance of my framework scales with the total number of facets in the scene.
-Furthermore, a specific performance limit was identified: the ability to maintain 30 fps, a common benchmark for interactive operation, is limited to a total facet count of approximately 500 (equivalent to about 10 objects).
-This measurement provides a performance baseline for evaluating the effects of future improvements, such as a GPU implementation.
+5.2.1 Convex Hull Mesh Generation Performance
+The measurement results are shown in Table 3.
+For point sets with up to 50 vertices, which correspond to the intended input of this implementation, the processing time was less than $1.7\text{\,}\mathrm{ms}$ in all tested dimensions within the measured range.
+In this implementation, convex hull generation is a processing stage whose computational cost varies depending on the input size within the architecture.
+To present baseline data for the architectural design regarding how convex hull generation time increases with input size, measurements were also conducted over the range of 100–1000 vertices as a scalability test beyond the intended input scale.
+5.2.2 Boolean Operation Performance
+The measurement results are shown in Table 4.
+In 3D space, for the combinations tested, the computation time was generally on the order of several tens of milliseconds.
+In contrast, in 4D space the computation time increased, reaching approximately $300\text{\,}\mathrm{ms}$ for Cube–Cube cases and exceeding $2\text{\,}\mathrm{s}$ for Random–Random cases in some instances.
+5.3 Rendering Performance
+The measurement results are shown in Table 5.
+Within the measured range, the FPS decreased as the total number of facets increased.
+Under the experimental conditions used in this study, when the total number of facets was 470 (corresponding to 10 objects), the average FPS was approximately 38.6.
+These values are recorded as reference values for the present implementation and environment.
+5.4 Practical Scope of Application and Usage Scenarios
+In this section, based on the measurement results of the 4D implementation, we describe the types of applications that can be realistically supported by the current implementation.
+Regarding visualization, in the present experiments, the frame rate decreased as the number of mesh facets increased.
+This may result in a loss of usability for continuous interaction.
+Therefore, when using this implementation for exploration or observation with continuous operations, a practical guideline is to keep the total number of facets at approximately 470 or fewer.
+For Boolean operations, immediate per-frame updates are not required, and a workflow of “operation → waiting for a certain period → result confirmation → next operation” can be viable.
+In the experiments, processing time tended to increase as complexity grew, and when it reached the order of several seconds, the pace of trial-and-error interaction was degraded.
+For this reason, the present implementation is realistically suited to research use cases or prototyping scenarios involving a small number of edits or validations, where waiting for results is acceptable.
+In contrast, usage scenarios that require continuous immediate responses, such as editing tasks that repeatedly apply Boolean operations at high frequency to large meshes with many facets, may become impractical with this implementation due to increased processing time as the number of facets grows.
+Accordingly, such scenarios are positioned outside the scope of the implementation and evaluation presented in this paper.
+However, as a workflow-level mitigation strategy, a proxy-based approach can be considered as one possible option.
+In this approach, trial-and-error operations are performed on simplified meshes, and the operations are applied to high-resolution models only at the final stage.
+5.5 Discussion and Limitations
+The definition of interactivity adopted in this paper holds regardless of computation time; however, practical usability is constrained as computation time increases.
+The experimental results suggest that when handling meshes with a large number of facets, processing time can reach the order of several seconds, leading to the existence of a practical transition region in which usage patterns that proceed by confirming results after each operation become constrained by waiting time.
+It should be emphasized that this transition region does not represent a strict threshold, but rather a practical consideration based on the experimental conditions of this study.
+Accordingly, this paper does not generalize acceptable response times or practical boundaries as a single numerical value.
+The present implementation employs approximate methods based on floating-point arithmetic, and intersection tests as well as inside-outside classification may become unstable depending on degenerate configurations and the choice of tolerances.
+Therefore, this paper explicitly states the limits of reliability in terms of these dependencies, and does not specify failure conditions as a single numerical boundary.
+The experiments in this implementation focus on computational performance and its scalability, and do not include qualitative evaluations such as user studies or structured feedback from domain researchers.
+In addition, the visualization-related limitations, namely information loss and dependence on sampling, are described in Section 3.6.1; quantitative evaluation of these limitations is beyond the scope of this paper.
 6 Conclusion and Future Work
-In this paper, I proposed an interactive framework for the unified simulation and visualization of $N$-dimensional space.
-The core of this framework lies in the design principle of separating topology and geometry.
-This allowed for the integration of functionalities such as Quickhull-based mesh generation, hyperplane slicing, and $N$-dimensional Boolean operations onto a single platform.
-The effectiveness of this design was demonstrated through the integration of an XPBD-based physics simulation.
-I also proposed a control scheme for high-dimensional exploration and the Plex data exchange format.
-The proposed algorithm can be extended to arbitrary dimensions; however, in this study, the evaluation was performed on the 4-dimensional implementation.
-Extension to five or more dimensions may introduce new performance bottlenecks due to the increased combinatorial complexity in intersection testing and tessellation.
-Furthermore, major future challenges include the development of a more robust $N$-dimensional non-rigid body model, application to new phenomena such as fluid simulations, and a GPU implementation of the parallelizable components currently handled by the CPU.
+In this paper, we proposed a unified architecture for visualization and simulation based on a design targeting $N$-dimensional space.
+We presented a configuration that integrates multiple processes, including convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing, into a single software architecture, and adopted the separation of topology and geometry as a design principle.
+In the 4D implementation, a non-rigid body simulation based on XPBD was implemented within the same architecture, and it was confirmed that updates to mesh vertex coordinates are reflected in cross-section generation and rendering.
+In addition, an interaction scheme for high-dimensional exploration was designed, and implemented in the 4D implementation.
+Furthermore, as one example of data representation for $N$-dimensional meshes, this study designed the Plex data format and defined its specification.
+Although the algorithms used in this study are described in a manner independent of the dimensionality $N$, performance evaluation and implementation-level verification are limited to four dimensions.
+Future work includes examining the impact of increased computational cost due to the growth of combinatorial complexity in intersection tests and simplicial subdivision for algorithms in five or more dimensions; investigating more numerically robust and physically accurate $N$-dimensional non-rigid physical models; exploring integration using other phenomena, such as fluid simulations, as input examples; and considering GPU implementations for components that are suitable for parallelization among the processes implemented in a CPU-based manner in this paper.
+In addition, conducting comparative evaluations with existing computational geometry libraries and previously reported methods, and quantitatively characterizing the trade-offs among computation time, memory usage, and robustness to degenerate configurations, also remain as future tasks.
+Furthermore, conducting qualitative evaluations also remains as future work.
+Such evaluations include collecting structured feedback from users and domain researchers on representative exploration and editing tasks.
+The design and the 4D implementation presented in this study may also serve as candidates for educational and entertainment applications aimed at understanding and manipulating high-dimensional phenomena.
 Acknowledgments
-I would like to express my sincere gratitude to Mr.
-Hiroaki Naito of Waseda Osaka High School for his constant and enthusiastic guidance throughout this research.
-I also wish to express my deepest appreciation to Professor Yasushi Homma of Waseda University for his valuable professional advice on the mathematical expressions related to Geometric Algebra used in this paper.
-Finally, I would like to thank the anonymous reviewers for their constructive comments, which helped to improve this manuscript.
+We would like to express our sincere gratitude to Mr.
+Hiroaki Naito of Waseda Osaka High School for his guidance throughout this research.
+We also wish to express our deepest appreciation to Professor Yasushi Homma of Waseda University for his valuable professional advice on the mathematical expressions related to geometric Algebra used in this paper.
+Finally, we would like to thank the anonymous reviewers for their constructive comments, which helped to improve this manuscript.

### v4

@@
 This paper proposes a unified software architecture for visualization and simulation based on a design targeting an $N$-dimensional space.
-The contribution of this work lies in presenting an architectural configuration that integrates multiple processes into a single software architecture: Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (including orientation and view transformations), and hyperplane slicing for visualization.
+The contributions of this study are twofold.
+First, it presents an architectural configuration that integrates multiple processes into a single software architecture: Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (pose transformation and view transformation), and hyperplane slicing for visualization.
+Second, it defines “Plex” (.plex) as a file format intended for the exchange of $N$-dimensional mesh data.
 The proposed approach adopts an approximate implementation that tolerates numerical errors and prioritizes implementation transparency over guarantees of numerical rigor.
@@
 However, examples remain limited that integrate Quickhull-based convex hull mesh generation, Boolean operations, coordinate transformations for high-dimensional exploration (orientation and view transformations), and hyperplane slicing for visualization into a single software architecture, under the assumption of an iterative workflow of user actions and result inspection.
-In this work, we propose a unified architecture intended for use by researchers and developers.
-While this architecture is designed for an $N$-dimensional space, the present study implements and evaluates it in 4D.
-The contribution of this work is to present an architectural configuration that integrates the above processing components as a single software architecture and to evaluate the 4D implementation experimentally.
+This study proposes a unified architecture intended for use by researchers and developers.
+When exchanging high-dimensional mesh data across projects, the availability of common formats is limited.
+Consequently, project-specific formats or extensions of existing formats are often employed.
+This study has two main contributions.
+First, it presents a configuration that integrates the above processing components into a single software architecture and evaluates a four-dimensional implementation through experiments.
+Second, it defines “Plex” (.plex) as a file format intended for the exchange of $N$-dimensional mesh data.
+The architecture is designed for $N$-dimensional space; however, the implementation and evaluation in this study are limited to 4D.
 In this paper, interactivity is defined as a workflow that presupposes repeated cycles of user actions and result inspection, in contrast to offline batch processing in which input conditions are fixed in advance and executed in a single run.
@@
 Cross-section generation based on this type of slicing plane has also been adopted in existing studies on interactive 4D visualization [5].
-In this implementation, no mechanism is provided to arbitrarily specify the orientation of the cutting hyperplane; cross-sections are always generated using the axis-aligned plane $w=w_{\text{slice}}$.
-When observing the cross-sections from arbitrary directions, changes in the viewing direction are handled as view transformations.
+Since rotating the object’s orientation via view transformation enables cross-sectional observation from arbitrary directions, this implementation does not employ a procedure to arbitrarily specify the orientation of the cutting hyperplane.
+Cross-sections are always generated using the axis-aligned plane $w=w_{\text{slice}}$.
 First, we determine on which side of the clipping hyperplane each of the facet’s four vertices is located.
@@
 In the 4D implementation, based on this design, a non-rigid body simulation based on XPBD is integrated, and it is confirmed that mesh vertex coordinates are updated over time and reflected in cross-section generation and rendering.
-However, evaluation of the validity or accuracy of the physical model is not included in the scope of this study.
+This serves as an implementation example demonstrating that new functionality can be integrated in the four-dimensional implementation without modifying the existing logic.
+Here, the role of XPBD is limited to serving as an example of a data source that provides dynamically changing geometry, and the validity or accuracy of the physical model is not within the scope of this study.
 4 Plex: An N-Dimensional Mesh Data Exchange Format
@@
 4.1 Background and Challenges
-In the $N$-dimensional simulation environments targeted by this study, there are cases where a means for exchanging high-dimensional mesh data between projects is required.
-However, for $N$-dimensional meshes, there are only a limited number of common exchange formats that are widely used across the research community.
-In practice, project-specific formats or extensions of existing formats are often employed.
+As described above, when exchanging high-dimensional mesh data across projects, the availability of common formats is limited.
 This situation can lead to additional effort when sharing data across different implementations or research groups, including data conversion, validation, and interpretation of specifications.
-To address this issue, this study defines a chunk-based file format, “Plex” (.plex), for the purpose of exchanging $N$-dimensional mesh data.
+Since the architecture proposed in this study is intended for use by researchers and developers, we define a chunk-based file format, “Plex” (.plex), to enable the exchange of $N$-dimensional mesh data.
 4.2 Relation to Existing Formats and Positioning of Plex
@@
 In contrast, in 4D space the computation time increased, reaching approximately $300\text{\,}\mathrm{ms}$ for Cube–Cube cases and exceeding $2\text{\,}\mathrm{s}$ for Random–Random cases in some instances.
+Numerical-error-induced holes or cracks in the mesh were not observed by visual inspection.
 5.3 Rendering Performance
@@
 In addition, the visualization-related limitations, namely information loss and dependence on sampling, are described in Section 3.6.1; quantitative evaluation of these limitations is beyond the scope of this paper.
-6 Conclusion and Future Work
+6 Conclusion
 In this paper, we proposed a unified architecture for visualization and simulation based on a design targeting $N$-dimensional space.
 We presented a configuration that integrates multiple processes, including convex hull mesh generation, Boolean operations, coordinate transformations, and hyperplane slicing, into a single software architecture, and adopted the separation of topology and geometry as a design principle.
+In addition, an interaction scheme for high-dimensional exploration was designed.
+Furthermore, as one example of data representation for $N$-dimensional meshes, this study designed the Plex data format and defined its specification.
 In the 4D implementation, a non-rigid body simulation based on XPBD was implemented within the same architecture, and it was confirmed that updates to mesh vertex coordinates are reflected in cross-section generation and rendering.
-In addition, an interaction scheme for high-dimensional exploration was designed, and implemented in the 4D implementation.
-Furthermore, as one example of data representation for $N$-dimensional meshes, this study designed the Plex data format and defined its specification.
+7 Future Work
 Although the algorithms used in this study are described in a manner independent of the dimensionality $N$, performance evaluation and implementation-level verification are limited to four dimensions.
-Future work includes examining the impact of increased computational cost due to the growth of combinatorial complexity in intersection tests and simplicial subdivision for algorithms in five or more dimensions; investigating more numerically robust and physically accurate $N$-dimensional non-rigid physical models; exploring integration using other phenomena, such as fluid simulations, as input examples; and considering GPU implementations for components that are suitable for parallelization among the processes implemented in a CPU-based manner in this paper.
+Future work includes examining the impact of increased computational cost due to the growth of combinatorial complexity in intersection tests and simplicial subdivision for algorithms in five or more dimensions; ensuring the topological correctness of meshes generated by Direct Quickhull; investigating more numerically robust and physically accurate $N$-dimensional non-rigid physical models; exploring integration using other phenomena, such as fluid simulations, as input examples; and considering GPU implementations for components that are suitable for parallelization among the processes implemented in a CPU-based manner in this paper.
 In addition, conducting comparative evaluations with existing computational geometry libraries and previously reported methods, and quantitatively characterizing the trade-offs among computation time, memory usage, and robustness to degenerate configurations, also remain as future tasks.